A Selected Bibliography of Publications by, and about, Samuel A. Goudsmit

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.06

Title word cross-reference

$3.50 [Bar30, Rid47]. ^{136} [Cha72]. ^{138} [Cha72]. ^{140} [Cha72]. ^{142} [Cha72]. g [Gou25g]. S^{a} [Ive10]. Z [LHLT64]. Z^{a} [Ive10].

-dependent [LHLT64]. -Werte [Gou25g].

1947 [Hen48, Whi48]. 1952 [Gou53b]. 1953 [Gou53a]. 1964 [Gou65a]. 1978 [Bed08b, Dre79]. 1983 [Moy84]. 1987 [Lug69]. 1988 [DGS89].

3d [Tho97].

A-Bomb [Rec91, LRD^{+}91, Bro93, Mac85]. Abides [Gla00]. Abraham [Bed08a, Bed08b]. Absorption [ZHG36]. Abstracts [GT66a, GT68, Gou69a]. Academy [Coc77, Jew77]. accept [Ano54a]. Acceptance [Gou72b]. Account [Lan54]. Ad [Gou72a]. Affirm [ACU^{+}54].

Again [GT66c, Gou73e, Rai85]. Age [Lan48, Sul78, Lan59a, Lan59b, Lap59]. AIP [Ano75]. Alamos [Bet91]. Alan [Gou78b]. Alfred [Hol93a, LH93]. Allies [Hol93b]. Alsos [Gou48c, Ano12, Gou47g, Gou47h, Gou47e, Gou48c, Gou51, Gou62, Gou83, Gou96, Pas69, Pas80, Ano48a, Gue50, Hen48, Moy84, Tre83, Whi48]. ALSOS. [Rid47]. Am [Gou74b]. amend [NG70]. American [EBU^{+}52, Gou47b]. Americans [Lan54]. Among [Tre83]. Analyses [BG32, BG68]. Analysis [Gou74a, Wer10]. Analyzed [Gou47c]. Ancient [Gou81]. Angeles [Moy84, Tre83]. Angeles/San [Tre83]. Angewandte [Gou50a]. Angle [Win89, Win87]. Angry [Gou63]. Angular [BL96, NLCS05]. Announcement [Gou58c, Gou58d, Gou58e, Gou68d]. Anomalies [GB33]. Anomalous [Ben38]. anonymity [WG67]. antiquities [RG82]. Application [Hei47a, Hei47b, MU56]. applications [Ike17]. Applied [Gou50a]. Appraisal [Hei49]. Arbeiten [Hei46]. arc [Cha72]. Archaeology [RG82]. Area [Ano48b]. Argon [GB27a, Hei27]. Arms [Pfe54]. Army [Gou47c]. Art [Gou72c, Gou81]. Arts [Gou58g]. Associate [Bar30]. Atom [Ano47, Ano49, Cas18, Gou47b, Gou47f, Gou49a, Gou49b, Hei49, GU26a, Las35, vCH18, GU26a]. atome [Las35]. Atomic [Ano54b, BG32, BG34, BG68, Gla00, Gou25d, GY30, Hei47a, Hei47b, Jun58, Lan48, Lan55, Lap59, Wal90, WHF^{+}96, Bet91, Gou47b, Gou47i, Gou48c, Hei46, HW49, Gou27, Lan59a, Lan59b]. atomique [Gou48c]. Atomization [Gou47a]. Atomkerne [HW49]. Atomkernenergie [Hei46]. Atoms [GR64, LHLT64]. atoombom [Hoe18, vC16]. Atoombouw [Gou25d]. Atoommodel [Gou27]. atualidade [LN54]. Auffassung [Gou21]. Ausnutzung [Hei46]. author [Gou72e, Gou75, PG61, WG67]. Await [Hei49]. Award [Ano54b, Ano54a]. Awards [Ano48b]. Awoke [Ano49].

Bacher [MU56]. Back [GB30, Hei27, Gaw91, GB29a, GB30, Vio72, Vio84]. Back-Effekt [GB30]. Backview [Gou81]. Balanced [Gou68b]. Barn [Gou59c]. Basel [Gou50a]. basic [Gou25f]. Be [Ano48b, Ano54a, Gou76b]. beams [TAL86]. becomes [Gla00]. Beeinflussung [Pau24]. before [BS85]. behavior [UG25]. Behind [Rid47]. below [Mac85]. Bernal [Gou53b]. Besprechung [Hei27]. Between [Gol89, Gou43]. Beyerchen [Gou78b]. Beyond [Cas09]. bezüglich [UG25]. Bias [Gou70a]. Big [Gou61b]. Biological [Gou39]. Birkhäuser [Gou50a]. Birth [BS85]. Bismuth [FG31a, FG31b, RHG52, BG28, GB27b, ZBG30]. Bleianzug [LR54]. body [BK59]. Bohr [Lak96]. Bomb [Ano47, Bro93, Gla00, Gou47b, Gou47f, Lau58a, WHF^{+}96, Cas09, GP92, Kar05, Mac85, vCH18, Rec91, Wal90, LRD^{+}91, Cas18, Hol93a, LH93]. Bombe [Kar05]. Book [And30, Ano48a, Bar30, Bro64, Cas18, Gou50a, Gou53a, Gou53b, Gou65a, Gue50, Hoe18, Hol93a, Lap59, Lug69, Moy84, Pfe54, Rid47, Tre83, Whi48, Wil17, Hei27]. Books [Gou78b, LH93]. Boston [Gou65a]. both [BL96]. bouw [Gou25a]. Brandeis [Gou58g]. Brighter [Jun58, Lau58a]. British [Ano48b]. Brookhaven [Gol89]. Brownian [UG29].

C [Tho97]. cadmium [Gou29a]. Cadmiumspektrum [Gou29a]. calcul [Las35]. Calculated [HUHM94]. Calculation [NLCS05, Las35]. calculations [KB98, TAL86]. California [Bar30]. Calmthout [Cas18, Hoe18, Wil17]. Came [Gou46a]. Carbon [Gou47a, GB27a, Hei27, Las35]. carbone [Las35]. Carlo [BS07, KIGT09, KB98, TAL86]. Case [Gou47c]. Catastrophe [LH93]. Catch [Tho77]. Ce [Cha72]. Century [FW60]. cerium [Cha72, Cha72]. CFSRM [HUHM94]. Change [GT69, Ano74]. Charge [Gou68d, Gou68a, Gou68e, Gou68f, Gou68g]. Charged [GS39]. Charges [Gou70b, Gou70c, Gou70d, Gou70e, Gou70f, Gou70g, Gou71c, Gou71d, Gou74d]. charts [Gou42]. Chemistry [Bar30, BS85]. Cherenkov [BL96]. Christopher [Gou65a]. Chronicles [Lan59a, Lan59b, Lap59]. chumba [LN54]. CII [RC25]. Circuit [GW43b]. Citation [Gou74a]. Claim [Gou49a, Gou65b]. Claim-Staking [Gou65b]. Clairborne [Lug69]. classically [Giu08]. clinical [TAL86]. Close [Gou46a, Lap59]. Close-Ups [Lap59]. code [KIGT09]. Codiscoverer [Sul78]. coefficients [Gou42]. coherent [HUHM94]. Collection [RG82, Ano11]. com [LN54]. Comments [GT74, Gou67a]. Committee [Gou72a]. Communication [Gou65c]. Communications [GY30, Gou74b]. Community [Gou78b]. Company [Bar30, Bro64]. comparison [Gou43]. complex [Gou25h]. componenten [GdLK25a]. Components [GdLK25c, GdLK25a, GdLK25b]. Compton [Ano75]. Computation [Put74]. computing [Vio72, Vio84]. concept [Ano54a, BS85]. Condensed [BS07]. Conference [DGS89, Lau58c]. Construction [Gou25d, Gou25a]. continued [DR92]. Conversations [Lan55]. Copenhagen [Fraxx]. corona [GW34]. Corp. [Ano54a]. Corporation [Ano54b]. Costs [Gou73b]. coupling [GU26a, GB27a, Hei27]. Credo [Gou63]. Crime [Gou53a]. Criteria [Gou72b]. Critical [GT72a]. Criticism [Gou72b]. Cross [ZHG36]. crossings [Vio72, Vio84]. Crystal [TWG48]. Cu [HUHM94]. Cumulative [Gou72e, Gou75, PG61]. Current [Lan54, LN54].

D [Gou53b, Gou78b]. damage [GS66a, GS66b]. Daniel [Lap59, Pfe54]. dans [Cha72, Gou22]. d’arc [Cha72]. Data [GT72a]. David [Gou76a, Gou76c]. Dead [Sul78]. debate [Bro93, DR92, GP92, Gou77b]. December [Bed08b]. Declassified [GP92]. découverte [Gou67b]. dedicate [Ano80]. defeat [Ols63]. deformed [BK59]. Dennison [Gou76a, Gou76c]. dependent [LHLT64]. Déplacement [Cha72, Cha72]. Derived [BG32, BG68]. describable [Giu08]. design [Wer10]. Determination [ZHG36]. Determinations [Gou25b, Gou25e]. d’étincelle [Cha72]. deutschen [Kar05]. Deutschland [Hei46]. Deutung [Pau24]. Develop [WHF^{+}96]. Development [Gou47b]. Did [LH93, Ols63]. Difficulties [Gou32]. Diffusion [BCGG36]. Digit [LG78, Rai85, Rai76]. Dirac [Lak96].

Discoverer [Ano54a]. Discoveries [Lau58c]. Discovery [Gou47f, Gou67b, Gou71e, Gou71a, Gou98, Sul78]. Displacement [Bre67, Cha72]. Disputed [Gou49a]. Disputes [Gou47b]. Distribution [Gou45, MYZ67, BL96, Put74, UG35]. distributions [KB98, NLCS05]. Doublet [Gou28, Gou21]. doublets [Gou22]. Down [Gou36]. Dr. [Ano49]. drama [Gla00]. Drive [Bro93]. Dubletts [Gou21]. durch [Pau24, UG25]. Dutch [Ano54b, Gou25a, Gou25c, Gou25d, GdLK25a, GU25, Gou25e, GU26b, Gou27, GCt67, Gou71a].

each [UG25]. Early [Lan48]. Eavesdropping [Hol93b]. echoes [GW43a]. Edited [Gou50a]. Editor [Gou58f, Gou59d, Gou61b, Gou63, LRD^{+}91]. Editorial [ATW79, Fel77, GT58a, Gou58a, GdMMP58, Gou58b, GT58b, GT58c, GP59a, Gou59a, GP59b, Gou59b, GP59c, Gou60a, Gou60b, Gou60c, Gou61a, Gou68c, Ano74]. Editors [Lug69]. Effect [GB29a, BG28, COGT69, Gaw91, Gou24a, GB30, LSH^{+}69, Vio72, Vio84]. Effekt [GB30]. effort [Bro93]. Efforts [Gla00]. Egyptian [RG82, Gou81]. Ehrenfest [MUG^{+}33]. einigen [GB27a, Hei27]. einiger [Pau24]. einzelnen [UG25]. elastic [KB98]. elastic-scattering [KB98]. electric [Ive10]. Electrodynamics [AK01]. Electron [Ano54a, Giu08, GG31, GR64, Sul78, Tho26, Wer10, GU26b, Gou67b, Gou71e, Gou71a, Gou98, KIGT09, KB98, LHLT64, Put74, TAL86, UG25, Gou67b]. electron-transport [TAL86]. electronenrotatie [Gou71a]. Electronic [Las35]. électroniques [Las35]. Electrons [Gou28, GS40c, GS40a, GS40b, UG26, UG67, UGB84, BS85, BL96, GS66a, GS66b, NLCS05]. Elektrons [UG25]. élémentaires [PG39]. Elementary [Gou77a, PG39]. Elements [WG33]. Eminent [EBU^{+}52]. Employment [Gou47c]. energies [Las35, Las35]. Energy [BG32, BG34, BG68, Hei47a, Hei47b, Seg85, GS66a, GS66b, Hei46, TAL86, UG35]. EPR [EES98]. Epsilon [Fra93a, Fra93b]. Equivalent [Gou28]. Errata [GT72b]. Ersetzung [UG25]. Essay [Ada08]. establishment [Ols63]. États [Las35]. Etc [Lan55]. Europe [DGS89]. European [Sul78]. experience [Gou58h]. Experiment [GT70c, Wer10]. Experiments [Gou39]. Explained [Lau58c]. Extension [Gou30a, Ike17].

Fact [Hen48]. Failed [WHF^{+}96]. failure [Gou47h]. Faith [ACU^{+}54]. Fallacies [Gou59d]. Fallout [Gou59d]. farewell [Lan53a, Lan53b]. Farm [Fra93a, Fra93b]. Feinstrukturen [GB27b]. Felder [Pau24]. Fermi [Gou59c]. Festschriften [Gou65d]. field [Gaw91]. field-optical [Gaw91]. fields [Pau24]. Fifty [Gou76b]. Figures [GF44, Gou81]. files [GP92]. findings [Bro93]. Fine [RC25, GB27b]. Finite [Win87, Win89]. Finite-Angle [Win89, Win87]. First [LG78, Rai85, Coc77, Rai76]. Flight [Gou48a, HRG51, HRG52]. following [Gou47b]. Footnote [Mac85]. force [UG25]. forces [BK59]. Forderung [UG25]. forgotten [Cas18]. Format [Gou68c]. forward [HUHM94]. Foundations [EES98]. Frage [Pau24]. Francisco [Tre83]. Frederic [Tho97]. Freedom [Gou50b]. French [Cha72, Gou48c, Gou67b, GC69, GC70a, Las35, PG39]. friends [Ano80]. Friendship [Gla00]. Front [Rid47]. Frontiers [Gou58g, Lau58c]. Fundamental [Gou73c, Gou73d].

Gallery [Pai00]. game [Gou78a]. Geant4 [KIGT09]. geheime [Kar05]. generalized [Ive10]. Genius [Pai00]. German [Hol93a, LH93, Ano47, BG28, Bro93, CG25, Gou21, Gou24d, Gou24a, Gou25f, GdLK25b, Gou25h, Gou25g, GU26a, GB27b, GB27a, Gou29a, GB30, Gou47h, Gou47b, Gou47c, Gou47d, Gou49a, GC67, GC70b, Hei27, Hei46, Hei49, HW49, Hol93b, Kar05, LH93, LRD^{+}91, Ols63, Pau24, Rec91, UG25, Wal90, Wal03, ZBG30]. Germany [Gla00, Gou46b, Gou46c, Gou47b, Gou48c, Gou48b, Hei46, Hei47a, Hei47b, WHF^{+}96]. Geschichte [Kar05]. Get [Ano48b]. Gibt [Gou29a]. Give [EBU^{+}52]. Gives [Ano54b]. got [Ols63]. Goudsmit [And30, Bar30, Bed08b, Bro64, Cas18, Dre79, Hei27, Hen48, Hoe18, Lak96, Lug69, Moy84, Rid47, Tre83, Whi48, ATW79, Ada08, Ano47, Ano48a, Ano49, Ano54a, Ano74, Ano75, Ano80, Ano11, Ano12, BS85, Bed08a, Gaw91, Gol89, Gol79, Gue50, HS09, Ike17, Ive10, KIGT09, Kae47, Kuh63a, Kuh63b, Kuh63c, LRD^{+}91, MU56, Put74, Rec91, Sul78, Tem09, Vio72, Vio84, Wal90, vC16, vCH18, Wil17]. Graduate [Gou58g]. Gray [Gou73a]. great [Ols63]. Grondniveau [Gou25c]. Ground [Gou25c]. group [GB27a, Hei27]. grouping [TAL86]. Growth [GT66b, GT67a]. Grundterm [Gou25f]. Gustav [Gou76d].

Hahn [Wal03, Wal03, Wal06]. Haigerloch [Mac85]. Hall [Fra93a, Fra93b]. happened [Ano54a, Gou69b]. Harvard [Lau58c]. Haughtiness [Tre83]. headed [Gou47b]. Heat [Gou47a]. Heaviest [WG33]. Heisenberg [Gou76f, Lak96, Bro93, Cas09, Gou47d, LRD^{+}91, Rec91, Wal90, Hol93a, LH93]. Held [Gou47f]. helium [GU25]. help [Ano80]. Henry [Rid47, Whi48]. Herewith [Ols63]. Hero [Tho97]. Hertz [Gou76d]. Hidden [Mac85]. higher [Gou25g]. higher-level [Gou25g]. Hill [Bar30, Bro64]. Hiroshima [Lap59, Ano49, Lan59a, Lan59b]. Historical [Seg85]. History [Hol93a, Jun58, Kar05, Kuh63a, Kuh63b, Kuh63c, LH93, Mac85, Moy84, BS07, KHFA67]. Hitler [Cas18, Gou78b, Hol93b, Kar05, vC16, vCH18, Hoe18]. Hitlers [Kar05]. Hoc [Gou72a]. höherer [Gou25g]. homem [LN54]. Honored [Ano48b, Ano54a, Ano54b]. House [Gou65a]. Houston [Gou30a]. Human [Gou81]. humor [HS09, Tem09]. hundred [Coc77]. Hunt [Cas18, vCH18]. Hydrogen [Lak96, GU25]. Hyperfeinstruktur [ZBG30, GB30]. Hyperfine [Bre67, FG31a, FG31b, Gaw91, GB29a, GB29b, Gou30b, Gou30c, GI31, Gou31, Gou32, GB33, Bel82, GB30, ZBG30]. Hypothese [UG25]. hypothesis [Ive10, UG25].

idea [Ano54a]. Iets [Gou25d]. ihrer [Pau24]. II [GS40c, Jew77, Kuh63b, Lan53b, LH93]. III [Kuh63c]. Illiterate [Gou74c].

Illustrated [Hol93a, LH93, Rid47]. illustrations [Bar30]. im [Gou24a, GU26a, Gou29a, LR54]. impedance [Gou42]. Important [Gou58c, Gou58d, Gou58e, Gou68d]. inauguration [Gou58g]. Incorporation [KIGT09]. index [Gou72e, Gou75, PG61]. Indexing [BBC^{+}66]. Indictment [Lau58a]. Inevitable [Gou47f]. influence [Pau24]. inner [UG25]. inneren [UG25]. Institute [Bar30]. Intelligence [Gou56, Rid47]. Intensität [CG25, GdLK25b]. intensiteiten [GdLK25a]. Intensities [GdLK25c, GdLK25a]. intensity [CG25, GdLK25b]. interaction [Gaw91]. Interested [Ano49]. International [Ano64, DGS89, Gou53b]. interned [Hol93b]. interpretation [Pau24]. Interscience [Gou53a]. Interviews [Kuh63a, Kuh63b, Kuh63c]. intrinsic [BS85]. Introduction [Kal85, Tre83]. inventory [KHFA67]. investigating [Gou47b]. Investigation [Gou53a]. Ionized [FG31a, FG31b, GI31, GR64]. Iron [Gou24b]. irrelevant [Gou77b]. Irving [Gou65a]. Isotope [Bre67, Gou29a]. Isotopenverschiebung [Gou29a]. isotopes [Cha72]. isotopic [Cha72, Cha72]. isotopique [Cha72]. Italy [DGS89]. IV [Ano12].

J [Gou53b]. jacht [Hoe18, vC16]. January [Gou50a, RG82]. jedes [UG25]. Jones [Tre83]. Journal [Gou68b, Gou50a]. July [Bed08b].

Kaempffert [Kae47]. Kelsey [RG82]. Kernmoment [BG28]. Kernwaffenversuche [Kar05]. Key [Ano54b]. Kirk [Gou53a]. Knopf [Hol93a, LH93]. Knowledge [Gou49b, Gou58g]. Kohlenstoffgruppe [GB27a, Hei27]. Komplexstruktur [Gou25h]. Koppelung [GB27a]. Kopplung [Hei27]. Kopplungsmöglichkeiten [GU26a].

L [Gou53a]. lab [Mac85]. L’Allemagne [Gou48c]. Lang [Lap59, Pfe54]. Lanthaan [Gou25e]. Lanthaan-spectrum [Gou25e]. Lanthanspectrum [Gou25a]. Lanthanspektrums [Gou24d]. Lanthanum [Gou25b, Gou25e, Gou24d, Gou25a]. laser [Gaw91]. Later [Gla00]. lateral [Ike17]. Law [Ben38, Gou28]. Lead [LN54, LR54, Lan54, Lan55, Pfe54, RHG52]. leaders [Ols63]. Legend [Hen48]. Length [GT67b, GT71]. Let [Fel77, Gou67c, Tho77]. Letter [DR92, Gou58f, Gou59d, Gou61b, Gou63, Gou78a, La 78, Tho77, Gou47b]. Letters [Gou65b, GT66a, GT67b, GT68, GT74, LRD^{+}91, Sau74, Ada08, GT67a, Gou72e, LG74, PG61]. level [Gou25g, Vio72, Vio84]. L’homme [GC69]. liaison [Las35]. Life [Lan59a, Lan59b, Lap59, Lug69]. light [BL96, BK59]. Line [And30, Bro64, PG30, Bar30, HUHM94]. Lines [GZ24, Gou45, CG25, Pau24, RC25]. link [Las35]. Linus [And30, Bar30, Bro64]. listening [Hol93b]. literature [Gou66c]. Lithium [GY30, GI31]. Lorentz [AK01]. loss [GS66a, GS66b]. Lost [Gou46b]. Low [WG33]. Lubricating [GO33].

M [Gou76a, Gou76c, Wil17]. Made [Lau58a]. Magnetic [GZ24, Gou33, HRG51, HRG52, BS85, Ive10, Pau24]. magnetische [Pau24]. magnetism [BS85]. Man [Gou63, LN54, LR54, Lan54, Lan55, Pfe54, Ano54a, GC69]. Manganese [Gou24c, Gou30b]. Männer [LR54]. Many [GG31, Sul78, LHLT64]. Many-Electron [GG31, LHLT64]. Martijn [Cas18, Hoe18]. Marwick [Ike17]. Marx [Gou53b]. Masers [Gou59e]. Mass [AK01, Gou48a, Gou65a, HRG51, HRG52, PG39]. masse [PG39]. Masses [RHG52]. Mathematics [Gou50a]. Mathematik [Gou50a]. May [RG82, Ano54a]. McGraw [Bar30, Bro64]. McGraw-Hill [Bar30, Bro64]. measure [GC70a]. measurement [GW43a]. Measurements [HRG51, HRG52]. Measuring [GS48]. mechanics [Gou29b]. Medal [Ano64, Ano75]. meeting [Gla00]. memorial [FW60]. Men [Lau58a, Pfe54]. mesure [GC70a]. Method [GS48, MU56, ZHG36]. methods [BS07]. Michigan [Bar30]. might [Gou76b]. Missiia [Gou62]. Mission [Rid47, Gou47b, Gou48c, Ano12, Pas69, Pas80]. Misunderstanding [SHW^{+}49]. Model [Gou27, GW43a, Ike17]. Modelers [Lak96]. Models [Lak96]. Modern [EES98, Moy84]. Moment [GY30, BG28, BS85]. Moments [Gou33]. monochromator [HUHM94]. Monte [KIGT09, BS07, KB98, TAL86]. Moon [Lan59a, Lan59b, Lap59]. Motion [Tho26, UG29]. Mr. [Kae47]. Multiple [GS39, GS40c, GS40a, GS40b, Ike17, KB98, NLCS05, Win87, Win89]. multiple-scattering [NLCS05]. Multiplet [GH28, Gou28, Gou30a]. Museum [RG82]. my [Gou69b]. Mystery [Gla00].

National [Coc77]. Nations [Gou47f]. Nature [GY30]. Nazi [Bro93, DR92, Gla00, GP92, Mac85, WHF^{+}96]. Nazis [Gou46a, Gou47i]. Nazism [Ols63]. near [GS66a, GS66b]. Neon [Gou25c, GB27a, Hei27, Gou25f]. Neonspectrum [Gou25c]. Neonspektrums [Gou25f]. Neutrons [BCGG36, Gou36, ZHG36]. Nine [Ano48b]. No [Gou50a, Hol93b]. Noise [GW43b, Gou43]. Nomenclature [Gou59f]. non [Giu08]. non-describable [Giu08]. Northwestern [Gou47b]. Note [GW34, Bel82, GU25]. Nuclear [GY30, Gou33, Gou47d, Hei53a, Hei53b, Kar05, MU56, Seg85, Bet91, Gou61c, BG28]. nuclei [BK59]. nucleus [HW49]. Number [Gou67e, UG35]. Numbers [Ben38, GF44, Gou78a].

O [LN54]. Obituary [Dre79, Gol79, Gou76c, Gou76d, Gou76e, MUG^{+}33]. Officer [Gou55]. Oils [GO33]. Old [Gol89, Gou63]. once [Ols63]. once-great [Ols63]. one [Hol93b]. online [Ano11]. Only [GT70c]. ontdekking [Gou71a]. op [Hoe18, vC16]. Operation [Fra93a, Fra93b]. Opmerking [GU25]. Oppenheimer [ACU^{+}54]. Optical [BG32, BG68, Gaw91]. Oral [Kuh63a, Kuh63b, Kuh63c]. Order [GR64, GB27b, LHLT64]. Ordering [GT66c, Gou24d]. Ordnung [Gou24d]. Other [Gou47f]. Otto [Wal03, Wal06]. Our [Gou48b]. Oversupply [GT70a]. Ovshinsky [COGT69, LSH^{+}69]. Oxford [Pfe54].

Page [Gou68d, Gou68a, Gou68e, Gou68f, Gou68g, Gou70b, Gou70c, Gou70d, Gou70e, Gou70f, Gou70g, Gou71c, Gou71d, Gou74d]. Page-Charge [Gou68d, Gou68a, Gou68e, Gou68g]. pages [Bar30, LH93]. panic [Fel77]. paper [Gou69b]. Papers [Ano12, Gou66b, Gou73c, Gou73d, BBC^{+}66, Gou58g]. Parity [La 78]. Particle [GS48]. Particles [GS39, PG39, UG35]. particules [PG39]. Paschen [GB30, Gaw91, GB29a, GB30]. past [RG82]. Pasternack [Gou76e]. Paul [Gou53a, Gou65a, MUG^{+}33]. Pauli [Bel82, FW60, Gou61c]. Pauling [And30, Bar30, Bro64]. Pay [Gou74d, Gou67c]. Percentages [DF79, MYZ67]. Personal [Jun58, Bet91]. Personified [Gou71b]. Perspective [Seg85, Bet91]. Ph.D [Bar30]. Phenomenon [LG78, Rai85]. photon [BS07]. Physical [DGS89, Gou72e, Gou75, PG61, Ada08, GT67a, Gou69a, Gou72e, Gou73b, LG74, PG61]. Physicists [Ano54b, Ano64, Tre83]. Physics [Bar30, Gla00, GY30, Gou46c, Gou47b, Gou50a, Gou58h, Gou71b, Gou78b, Hei53a, Hei53b, Lau58c, Moy84, BBC^{+}66, Cas09, FW60, KHFA67, RG76, HW49]. Physik [Gou50a, HW49]. pioneer [Cas18]. Pioneers [Ano54a]. Pitfalls [Gou77a]. Places [Gou65a]. Plane [Gou45]. point [Kro60]. policies [LG74]. Policy [EBU^{+}52, GT69]. Politics [Gou78b]. Portrait [Pai00]. Portuguese [LN54]. Positive [PBG55]. positrons [NLCS05]. possibilities [GU26a]. Possible [GW43a]. Postscript [Fraxx]. Power [Gou47d, Bet91]. Powers [Hol93a, LH93]. pp [Gou53a, Gou53b, Gou65a, Hen48, Hol93a, Lap59, Lug69, Moy84, Pfe54, Rid47, Whi48]. pp. [Tre83]. Predictions [Gou30b]. Prejudice [RG76]. prematurely [Fel77]. Present [Gou32]. Press [Pfe54]. Prevent [LH93, SHW^{+}49]. Preview [Gou47e]. Price [Bar30]. Pride [Tre83]. Prime [Gou67e]. Probability [Gou77a]. Problem [Gou68d, Gou68a, Gou68e, Gou68f, Gou68g, UG29, Rai76]. problemas [LN54]. Problems [Gou65c, Lan54, BS07, LN54]. Proceedings [DGS89]. produced [BL96]. Professor [Bar30, Gou47b]. Profiles [Lan53a, Lan53b]. Profound [Bet91]. Program [PBG55, Vio72, Vio84]. Progress [Gou49a, Hei49, Gou47b]. Project [Ano47, Gou47d]. Prometheus [Cas18]. Proper [Gou65a, Gou65a]. propose [Ano54a]. Proposed [GS48]. Publication [Gou73b]. Publicity [Gou60d]. Publishers [Gou53a, Gou53b]. Publishing [Gou65a].

Quantenvektoren [GU26a, GB27a, Hei27]. Quantum [Gou29b, Cas09, Cas18, GU26a, GB27a, Hei27, KHFA67]. question [Pau24].

R [Gou50a, Tre83]. Race [Gou46b]. radar [GW43a]. radiation [GS66a, GS66b]. Radioactivity [Gou39]. Random [Gou45]. Range [Ike17]. ray [CG25]. Reactions [Lan54]. read [Gou58g]. ready [Ano54a]. Receive [Ano64]. receives [Ano75]. Recollections [HS09, Tem09]. Rectifiers [TWG48]. Reduction [Gou71c, Gou71d]. Referee [Gou66a, GT70b, LG74]. Refereeing [Gou66b, Gou72b, NG70]. referees [Gou67c]. Reflection [Gou42]. Refugee [Seg85]. Regarding [Gou73c, Gou73d]. Reich [Gou78b]. Related [GY30]. relatif [Cha72]. Relations [BG34, Gou30a]. Relative [Cha72]. Relatives [Ano80]. Relativistic [Gou21]. Relativistische [Gou21]. Relaxation [GT70c]. Release [Hei49]. Remark [Gou30c]. Renormalization [AK01]. reopen [Bro93, GP92]. Replacement [UG25]. Replies [Kae47]. Reply [Lau58b]. report [KHFA67]. Reports [GT70b, Hei49]. representation [KB98]. Repression [Wal06, Wal03]. Requirement [GT72b, UG25]. Research [Ano54a, Gou47b, Gou49a, Hei47a, Hei47b, Hei49, Ano54a, Ano54b]. Residents [Ano48b]. Resolution [GZ24]. resonance [HUHM94]. Responsibility [Wal06, Wal03]. Rest [Ano93]. Restriction [GT67b]. Restructuring [DGS89]. retires [Ano74]. retrieving [Gou66c]. Review [Ada08, And30, Ano48a, Bar30, Bro64, Gou50a, Gou53a, Gou53b, Gou65a, GT67a, Gou72e, Gue50, Hei27, Hoe18, Hol93a, Lap59, Lau58c, LG74, Lug69, Moy84, PG61, Pfe54, Rid47, Tre83, Whi48, Wil17, Cas18, Hen48, Ols63, Gou69a, Gou72e, Gou75, PG61]. Reviewer [WG67]. Reviews [Gou73b]. Revised [GT66c]. revisited [Win89]. Richards [Gou65a]. Risks [Gou61b]. road [Bet91]. Robert [Lug69]. Roentgen [Gou28]. role [Bro93]. Rome [DGS89]. Röntgenspektrallinien [CG25]. room [Ano80]. rotating [GU26b]. roteerende [GU26b]. Rounded [MYZ67]. Rounding [DF79]. roupa [LN54]. Rules [GG31, GT66b]. Russia [Ano49]. Rutherford [Lak96].

S [Hei27]. S. [Hei27]. Saboteur [Bro93]. Said [Hei49, Hol93b, Kle09]. Sallhofer [Lak96]. Sam [Ada08, Hoe18, Kle09, vC16, vCH18, Cas18, Wil17]. Samuel [Ano48a, Bar30, Bro64, Gue50, Hen48, Lug69, Moy84, Rid47, Tre83, Whi48, ATW79, Ano12, Bed08a, Bed08b, Dre79, Gol79, HS09, Kae47, Kuh63a, Kuh63b, Kuh63c, RG82, Sul78, Tem09]. San [Tre83]. Sänger [Gou50a]. Sapienza [DGS89]. satellite [Pau24]. Satelliten [Pau24]. Saunderson [Ike17, KIGT09, Put74]. Savant [Bro93]. Says [Ano49, Ano54a]. Scandium [GZ24, Gou24a]. Scandiumspektrum [Gou24a]. Scared [Gou74b, Sau74]. scattered [Ols63]. Scattering [GS66a, GS66b, GS39, GS40c, GS40a, GS40b, GS48, HUHM94, Ike17, KB98, NLCS05, Win87, Win89]. scholars [Hol93b]. School [Gou58g]. Schrödinger [Lak96]. Schuman [Hen48, Rid47, Whi48]. Schuster [Lap59]. Science [Ano64, Gou47f, Gou50b, Lau58c, Pai00, Bet91, DR92, Gou46d, Gou47h, Gou53b]. Sciences [DGS89, Gou58g, Coc77]. Scientific [Gou56, Lan54, Rid47, Ols63]. Scientist [Ano54b, Gou47b, Gou49a, LH93, Ols63, RG82]. Scientists [ACU^{+}54, EBU^{+}52, Gou47c, Hol93b, Jun58, Lan55, Lap59, Seg85, Gou78b]. Scribe [Gou74c]. sealing [Lan53a, Lan53b]. Second [LHLT64, LSH^{+}69]. Second-order [LHLT64]. Secrecy [Gou46d, Gou73e]. Secret [Gou47f, Hol93a, Kar05, LH93, Gou48c]. secrets [Gou47i]. Sections [ZHG36]. Security [Gou55, PBG55]. seems [Mor07]. Selecting [Gou66a]. Selection [GG31]. Separations [GH28, Gou28, GB29b, Gou31]. September [DGS89]. Sequences [Gou67e]. Series [Ano12]. Session [Kuh63a, Kuh63b, Kuh63c]. Shells [GR64]. shift [Gou29a]. Sigmund [Ike17]. signal [Gou43]. Significant [GF44]. Silence [Ano93]. silicon [GS66a, GS66b]. Simon [Lap59, Gou76e]. Slater [Gou30a]. Slow [BCGG36]. Slowing [Gou36]. Small [Gou61b, Mac85, UG35]. Sobriety [Tre83]. some [GB27a, Hei27, Ols63, Pau24]. Something [Gou25d]. Sommerfeld [Lak96]. Song [Gou74e]. Sources [KHFA67]. Soviet [Gou49b]. sparkle [Cha72]. Special [GT72b]. Spectra [BG32, BG68, Gou25d, MU56, PG30, UG26, UG67, UGB84, Cha72, Gou25h, GU25, Gou25g, GU26b, Gou27, GB27a, Hei27, And30, Bro64, Bar30]. spectral [CG25, Pau24]. spectres [Cha72, Gou22]. Spectrometer [Gou48a, HRG51, HRG52]. Spectrum [Gou24b, Gou24c, Gou25b, Gou25c, Gou24d, Gou24a, Gou25a, Gou25f, Gou25e, GB27b, Gou29a, GW34, RC25, Wer10]. Spektrallinien [Pau24]. Spektren [Gou25g, GB27a, Hei27, Gou25h]. Spin [Ano54a, AK01, Mor07, Sul78, Giu08, Gou61c, Gou67b, Gou71e, Gou71a, Gou76b, Gou98]. Spinning [Tho26, UG26, UG67, UGB84]. spins [Ive10]. spread [Ike17]. spy [Cas18]. staffs [Ano74]. stage [Gla00]. Staking [Gou65b]. State [Gou25c]. Statement [Gou47b]. States [BG32, BG68, DGS89, WG33, Las35]. Statistical [GF44, UG35]. Statistics [GW43b]. stopping [BL96]. Strategic [La 78]. string [Lan53a, Lan53b]. strong [Gaw91]. Structure [And30, Bre67, Bro64, FG31a, FG31b, Gou27, GB29a, GB29b, Gou30b, Gou30c, GI31, Gou31, Gou32, GB33, PG30, UG26, UG67, UGB84, Bar30, Bel82, Gou25h, GU26b, GB27b, GB30, ZBG30]. structures [HUHM94, RC25]. structuur [GU26b, Gou27]. Stufe [Gou25g]. Subjects [GY30]. Suit [LN54, LR54, Lan54, Lan55, Pfe54]. Sums [MYZ67]. Suns [Jun58, Lau58a]. Superficiality [Gou67d]. Swan [Gou74e]. switching [Ike17]. Switzerland [Gou50a]. Symbols [Gou37]. Symmetry [Gou37].

T [Gou65a]. Tables [GF44]. Tales [Lan48]. targets [GW43a]. Task [Gou48b, Gou55]. TE [Cha72]. Technical [Hei47a, Hei47b, Hei46]. Technischen [Hei46]. Technology [Bar30]. temps [GC69, GC70a]. term [Gou25f, GB27b]. Termbepalingen [Gou25e]. Terme [Gou25g]. Termordnung [GB27b]. Terms [Gou25b, Gou25e, Gou25g]. Tests [Kar05]. Thales [Lak96]. Their [EBU^{+}52, Pau24]. them [Ols63]. Theoretical [Bar30, FW60, Pau24]. theoretischen [Pau24]. Theories [Gou73c, Gou73d]. Theory [Ano54b, Gou31, Gou32, Ike17, KIGT09, LHLT64]. there [Gou29a, Ols63]. Thermal [ZHG36]. Thick [LN54, LR54, Lan54, Lan55, Pfe54]. Third [Gou78b]. Thomas [Hol93a, LH93]. Thomson [Lak96]. thought [Hol93b]. Thousand [Jun58, Lau58a]. threshold [GS66a, GS66b]. Ties [Sul78]. tijd [GCt67]. Time [Gou48a, GCt66, HRG51, HRG52, Lug69, GC69, GC70a, GCt67, GC67, GC70b, Lug69]. Time-of-Flight [Gou48a, HRG51, HRG52]. Times [LH93]. Today [Ano48b]. Tomash [Moy84, Tre83]. Too [Sau74]. Town [Mac85]. Trackdown [Ols63]. Tracks [GS48]. Trade [Gou72c]. transcripts [Fra93a, Fra93b]. transmission [HUHM94]. transport [BS07, KIGT09, Put74, TAL86, Wer10]. treatment [Ike17]. Troubles [GT71]. turned [Cas18]. turning [Kro60]. Twentieth [FW60]. Two [BK59, GT72b, Giu08]. two-valuedness [Giu08].

U.S. [Ano64]. Uhlenbeck [BS85, Ive10]. Uncertain [Hol93a]. uncertainty [Cas09]. United [DGS89]. Università [DGS89]. University [Bar30, Pfe54, Gou47b, Gou58g]. unknown [GdLK25d, Gou72d, Lor36]. Unmasks [Ano47]. unmechanical [UG25]. unmechanischen [UG25]. Until [Ano49]. Unusual [Gou67e]. Ups [Lap59]. use [GW43a, Hei46]. using [TAL86].

V [Tre83]. Value [Gou66b]. valuedness [Giu08]. values [Gou25g]. Various [Lan54]. vectors [GU26a, GB27a, Hei27]. Verantwortung [Wal03]. Verdrängung [Wal03]. Verhaltens [UG25]. Verlag [Gou50a]. versus [Gaw91]. view [Gou21]. viewpoints [Bet91]. Views [EBU^{+}52, Lan54, RG82]. Visa [EBU^{+}52]. Viscosity [GO33]. visibles [Gou22]. Vision [Ada08]. vo [Bar30]. Vol [Gou50a]. volume [FW60, Moy84]. volumes [Gou72e, Gou75, PG61]. Voluntary [Gou74f]. vom [UG25]. vs [Hen48].

War [Gou46c, Gou49a, Hol93a, Jew77, LH93, Tho97, Bro93]. Wardenburg [Tho97]. Warning [Gou59g]. wartime [Cas18]. Was [Ano93, Bro93, Hol93a, Ano54a, Hol93b]. water [BL96]. waterstof [GU25]. wax [Lan53a, Lan53b]. weapons [Bet91]. well [Gou76b]. went [Ols63]. Werner [Gou76f]. Werte [Gou25g]. where [Ols63]. Who [Lau58a, Ano54a, Cas18]. Wide [Gou73a]. Will [Ano48b]. winner [Ano54a]. Wismut [BG28]. Wismuts [ZBG30]. Wismutspektrums [GB27b]. Wolfgang [FW60]. Words [Gou65a]. Work [Ano54b, Lap59, BS85, Hei46]. world [Ano54a, Jew77, LH93]. worth [Gou66c]. writer [Gou47b].

X [CG25]. X-ray [CG25]. xiv [Hen48, Rid47].

Years [Gla00, Coc77, Gou76b]. York [Bar30, Gou53a, Gou53b, Hen48, Hol93a, Lap59, Lug69, Pfe54, Rid47, Whi48].

Zeeman [BG28, Gou24a, GdLK25c, GdLK25a, GdLK25b]. Zeeman-componenten [GdLK25a]. Zeemaneffekt [BG28, Gou24a]. Zeemankomponenten [GdLK25b]. Zeit [GC67, GC70b]. Zeitschrift [Gou50a]. Zijn [Hoe18]. zur [Hei46, Gou24d, Pau24, ZBG30]. Zwang [UG25].

References

Allison:1954:SAF

[ACU^{+}54] Samuel K. Allison, Edward Condon, Harold C. Urey, Farrington Daniels, F. W. Loomis, Edward Shils, Linus Pauling, S. A. Goudsmit, Julian Schwinger, Albert Einstein, Hugh C. Wolfe, Cyril S. Smith, Victor F. Weisskopf, and Hans Bethe. Scientists affirm faith in Oppenheimer. Bulletin of the Atomic Scientists, 10(5):188–190, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adair:2008:EPR

[Ada08] Robert K. Adair. Essay: Physical Review Letters; Sam Goudsmit’s vision. Physical Review Letters, 100(2):020001:1–020001:4, January 18, 2008. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Appel:2001:MSR

[AK01] Walter Appel and Michael K.-H. Kiessling. Mass and spin renormalization in Lorentz electrodynamics. Annals of Physics, 289 (1):24–83, 2001. CODEN APNYA6. ISSN 0003-4916 (print),

1096-035x (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0003491600961190.

Andrews:1930:BRS

[And30] Donald H. Andrews. Book review: The Structure of Line Spectra (Pauling, Linus; Goudsmit). Journal of Chemical Education, 7(8): 1978, August 1930. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Anonymous:1947:GUG

[Ano47] Anonymous. Goudsmit unmasks German atom bomb project. Chemical & Engineering News, 25(44):3216–3217, November 1947. CODEN CENEAR. ISSN 0009-2347 (print), 1520-605X (electronic).

Anonymous:1948:BRA

[Ano48a] Anonymous. Book review: Alsos, by Samuel A. Goudsmit. The Times Literary Supplement, ??(??):??, ???? 1948.

Anonymous:1948:WGB

[Ano48b] Anonymous. Will get British awards: Nine residents of area are to be honored today. New York Times, ??(??):27, September

7, 1948. URL http://search.proquest.com/hnpnewyorktimes/
docview/108372530/.

Anonymous:1949:HAR

[Ano49] Anonymous. Hiroshima awoke Russia: Not interested in atom until then, says Dr. Goudsmit. New York Times, ??(??):5, December 8, 1949. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/105889902/.

Anonymous:1954:RES

[Ano54a] Anonymous. Research: Electron spin pioneers honored: “discoverer” of a concept may be the man who happened to propose the idea when the world was ready to accept it, says Research Corp. award winner Goudsmit. Chemical & Engineering News, 32(6): 480–486, February 1954. CODEN CENEAR. ISSN 0009-2347 (print), 1520-605X (electronic).

Anonymous:1954:SWH

[Ano54b] Anonymous. Scientists’ work of 1925 honored: Research Corporation gives award to 2 Dutch physicists for key to atomic theory. New York Times, ??(??):13, January 23, 1954. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/112966315/.

Anonymous:1964:UPR

[Ano64] Anonymous. 2 U.S. physicists to receive international science medal. New York Times, ??(??):12, June 17, 1964. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/115824554/.

Anonymous:1974:GRE

[Ano74] Anonymous. Goudsmit retires; editorial staffs change. Physics Today, 27(12):75, December 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.3129054.

Anonymous:1975:GRA

[Ano75] Anonymous. Goudsmit receives AIP Compton Medal. Physics Today, 28(1):105, January 1, 1975. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.3068786.

Anonymous:1980:RFH

[Ano80] Anonymous. Relatives and friends help dedicate Goudsmit room. Physics Today, 33(1):77, January 1, 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.2913910.

Anonymous:1993:RWS

[Ano93] Anonymous. The rest was silence. New York Times, ??(??): BR3, March 14, 1993. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/109207251/.

Anonymous:2011:GCO

[Ano11] Anonymous. Goudsmit collection online. Physics Today, 64(9): 34, September 1, 2011. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://physicstoday.

scitation.org/doi/full/10.1063/PT.3.1256.

Anonymous:2012:SGP

[Ano12] Anonymous. Samuel A. Goudsmit papers, 1921–1979: Series IV. Alsos Mission. ALSOS Digital Library for Nuclear Issues Web

site, 2012. URL http://alsos.wlu.edu/information.aspx?id=
4210.

Adair:1979:ESG

[ATW79] Robert K. Adair, George L. Trigg, and Gene L. Wells. Editorial: Samuel A. Goudsmit (1902–1978). Physical Review Letters, 42(1): 1–2, January 1, 1979. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Barnes:1930:BRS

[Bar30] James Barnes. Book review: The structure of line spectra: By Linus Pauling, Ph.D., Associate Professor of Theoretical Chemistry, California Institute of Technology and Samuel Goudsmit, Ph.D., Associate Professor of Theoretical Physics, University of Michigan. 258 pages, illustrations, 8 vo. New York, McGraw-Hill Book Company, Inc., 1930. Price $3.50. Journal of The Franklin Institute, 210(3):393, 1930. CODEN JFINAB. ISSN 0016-0032 (print),

1879-2693 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0016003230907004.

Blizard:1966:IPP

[BBC^{+}66] E. P. Blizard, F. G. Brickwedde, J. H. Crawford, S. A. Goudsmit, A. Herschman, S. Pasternack, A. B. Smith, and K. Way. Indexing physics papers. Physics Today, 19(1):54–??, January 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL https://physicstoday.scitation.org/doi/
abs/10.1063/1.3047979.

Bayley:1936:DSN

[BCGG36] D. S. Bayley, B. R. Curtis, E. R. Gaerttner, and S. Goudsmit. Diffusion of slow neutrons. Physical Review (2), 50(5):461–463, September 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Bederson:2008:SAGa

[Bed08a] Benjamin Bederson. Samuel Abraham Goudsmit (1902–1978). Physical Review Letters, 101(1):10002:1–10002:5, July 4, 2008. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Bederson:2008:SAGb

[Bed08b] Benjamin Bederson. Samuel Abraham Goudsmit: July 11, 1902– December 4, 1978. National Academy of Sciences, Washington,

DC, USA, 2008. 29 pp. URL http://www.nasonline.org/
publications/biographical-memoirs/memoir-pdfs/goudsmit-
samuel.pdf.

Belloni:1982:PNH

[Bel82] Lanfranco Belloni. Pauli’s 1924 note on hyperfine structure. American Journal of Physics, 50(5):461–464, May 1982. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Benford:1938:LAN

[Ben38] Frank Benford. The law of anomalous numbers. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 78(4):551–572, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL http://links.jstor.org/sici?sici=0003-049X%
2819380331%2978:4%3C551:TLOAN%3E2.0.CO%3B2-G. See com-

ments about Benford’s biased rounding practices [DF79].

Bethe:1991:RAP

[Bet91] Hans Albrecht Bethe. The road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science. Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1991. ISBN 0-88318-707-8, 0-671-74012-1. xvii + 286 pp. LCCN U264 .B455 1991. US$24.95.

Back:1928:KZW

[BG28] E. Back and S. Goudsmit. Kernmoment und Zeemaneffekt von Wismut. (German) [Nuclear moment and Zeeman effect in bismuth]. Zeitschrift für Physik, 47(3–4):174–183, February 1928. CODEN ZEPYAA. ISSN 0044-3328. URL http://link.

springer.com/article/10.1007/BF02055794;   http://www.
springerlink.com/content/rnwt7303k7677336/.

Bacher:1932:AES

[BG32] Robert F. (Robert Fox) Bacher and Samuel Goudsmit. Atomic Energy States, as Derived from the Analyses of Optical Spectra. McGraw-Hill, New York, NY, USA, 1932. xiii + 562 pp.

Bacher:1934:AER

[BG34] R. F. Bacher and S. Goudsmit. Atomic energy relations. I. Physical Review (2), 46(11):948–969, December 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Bacher:1968:AES

[BG68] Robert F. (Robert Fox) Bacher and Samuel Goudsmit. Atomic Energy States, as Derived from the Analyses of Optical Spectra. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1968. xiii + 562 pp.

Brink:1959:TBF

[BK59] D. M. Brink and A. K. Kerman. Two body forces in light deformed nuclei. Nuclear Physics, 12(4):314–326, 1959. CO-DEN NUPHA7. ISSN 0029-5582 (print), 1873-1546 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0029558259900306.

Bowler:1996:ADC

[BL96] M. G. Bowler and M. D. Lay. Angular distribution of Cherenkov light from electrons both produced and stopping in water. Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment, 378 (3):468–471, 1996. CODEN NIMAER. ISSN 0168-9002 (print),

1872-9576 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0168900296005451.

Breit:1967:IDH

[Bre67] G. Breit. The isotope displacement in hyperfine structure. In W. R. Hindmarsh, editor, Atomic Spectra, The Commonwealth and International Library: Selected Readings in Physics, pages 289–298. Pergamon Press, Oxford, UK and New York, NY, USA, 1967. ISBN 0-08-012103-9. URL http://www.sciencedirect.

com/science/article/pii/B9780080121031500256. Reprinted

from Physical Review 42, 348 (1932).

Bronowski:1964:BRS

[Bro64] J. Bronowski. Book review: The Structure of Line Spectra, by Linus Pauling and Samuel Goudsmit. McGraw-Hill Book Company. Scientific American, 210(6):136, June 1964. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v210/n6/pdf/scientificamerican0664-130.pdf.

Broad:1993:NFR

[Bro93] William J. Broad. New findings reopen debate on Heisenberg’s role in German war effort.: Was Heisenberg saboteur or savant of Nazi drive for A-bomb? New York Times, ??(??):C1– C2, September 1, 1993. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/109030191/.

Baracca:1985:BCI

[BS85] Angelo Baracca and Edvige Schettino. Birth of the concept of intrinsic magnetic moment of electrons in chemistry and magnetism before the work of Uhlenbeck and Goudsmit (1915–1921). Rendiconti della Accademia Nazionale delle Scienze detta dei XL. Memorie di Scienze Fisiche e Naturali. Serie V. Parte II, 9:263–267, 1985. ISSN 0392-4130. Proceedings of the fifth national congress on the history of physics (Italian) (Rome, 1984).

Bhan:2007:CHM

[BS07] Katherine Bhan and Jerome Spanier. Condensed history Monte Carlo methods for photon transport problems. Journal of Computational Physics, 225(2):1673–1694, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999107000770.

Cassidy:2009:BUH

[Cas09] David C. Cassidy. Beyond uncertainty: Heisenberg, quantum physics, and the bomb. Bellevue Literary Press, New York, NY, USA, 2009. ISBN 1-934137-13-8. 480 pp. LCCN QC16.W518 C37 2008.

Castelvecchi:2018:FQP

[Cas18] Davide Castelvecchi. The forgotten quantum pioneer who turned wartime spy [book review of Sam Goudsmit and the Hunt for Hitler’s Atom Bomb, by Martijn van Calmthout, Prometheus (2018)]. Nature, 563(7731):320–321, November 15, 2018. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://www.nature.com/articles/d41586-018-07267-3.

Coster:1925:IRG

[CG25] D. Coster and S. Goudsmit. Über die Intensität der Röntgenspektrallinien. (German) [On the intensity of X-ray spectral lines]. Die Naturwissenschaften, 13(1):11–12, January 1925. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://

link.springer.com/article/10.1007/BF01558797;   https:
//repository.aip.org/islandora/object/nbla%3A242705#page/
1.

Champeau:1972:DIS

[Cha72] R.-J Champeau. Déplacement isotopique dans les spectres d’arc et d’étincelle du cérium: I. Déplacement isotopique relatif des isotopes ^{136}Ce, ^{138}Ce, ^{140}Ce TE ^{142}Ce. (French) [Isotopic displacement in arc and sparkle spectra of cerium: I. Relative isotopic displacement of isotopes ^{136}Ce, ^{138}Ce, ^{140}Ce TE ^{142}Ce]. Physica, 62(2):209–224, 1972. CODEN PHYSAG. ISSN 0031-8914 (print),

1873-1767 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0031891472901243.

Cochrane:1977:NAS

[Coc77] Rexmond C. (Rexmond Canning) Cochrane. The National Academy of Sciences: the first hundred years, 1863–1963. National Academy Press, Washington, DC, USA, 1977. ISBN 0-309-02518-4. xv + 694 pp. LCCN Q11.N2862 C6.

Chopra:1969:OE

[COGT69] Kasturi L. Chopra, Stanford R. Ovshinsky, S. A. Goudsmit, and George L. Trigg. The Ovshinsky effect. Physics Today, 22(3):9–??, March 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL https://physicstoday.scitation.org/
doi/abs/10.1063/1.3035495.

Diaconis:1979:RP

[DF79] Persi Diaconis and David Freedman. On rounding percentages. Journal of the American Statistical Association, 74(366):359–364, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284288. The

authors extend prior work on correctness of sums of rounded percentages [MYZ67], and criticize biased rounding practices in [Ben38].

DeMaria:1989:PIC

[DGS89] Michelangelo De Maria, Mario Grilli, and Fabio Sebastiani, editors. Proceedings of the International Conference on the Restructuring of Physical Sciences in Europe and the United States, 1945– 1960: Università “La Sapienza,” Rome, Italy, 19–23 September 1988. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1989. ISBN 9971-5-0740-4. LCCN QC9.U5 I57 1988.

Dietz:1992:LNS

[DR92] Chris Dietz and Ernest A. Ryavec. Letter: “Nazi science” debate continued. Bulletin of the Atomic Scientists, 48(9):46, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [GP92].

Dresden:1979:OSG

[Dre79] Max Dresden. Obituary: Samuel A. Goudsmit, 1902–1978. Nature, 282(5741):889–890, December 20, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://

www.nature.com/articles/282889a0.

Einstein:1952:ESG

[EBU^{+}52] Albert Einstein, Hans A. Bethe, Harold C. Urey, James Franck, Samuel Goudsmit, Cyril S. Smith, Arthur H. Compton, and William P. Murphy. Eminent scientists give their views on American visa policy. Bulletin of the Atomic Scientists, 8(7):217–220, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Eaton:1998:FME

[EES98] Gareth R. Eaton, Sandra S. Eaton, and Kev Minullinovich Salikhov. Foundations of Modern EPR. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1998. ISBN 981-02-3295-0, 981-281-676-3. xiii + 818 pp.

LCCN QC763 .E28 1998. URL http://www.worldscientific.
com/worldscibooks/10.1142/3624\#t=toc.

Feld:1977:ELP

[Fel77] Bernard T. Feld. Editorial: Let’s not panic prematurely. Bulletin of the Atomic Scientists, 33(5):8–9, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tho77].

Fisher:1931:HSIb

[FG31a] Russell A. Fisher and S. Goudsmit. Hyperfine structure in ionized bismuth. Physical Review (2), 37(9):1057–1068, May 1931. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Fisher:1931:HSIc

[FG31b] Russell A. Fisher and S. Goudsmit. Hyperfine structure in ionized bismuth. Physical Review (2), 37(12):1674, June 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Frank:1993:OEFa

[Fra93a] Sir Charles Frank, editor. Operation Epsilon: the Farm Hall transcripts. Institute of Physics Publishers, Bristol, UK, 1993. ISBN 0-7503-0274-7. ix + 313 pp. LCCN Q141 .O56 1993. US$15.00.

Frank:1993:OEFb

[Fra93b] Sir Charles Frank, OBE, FRS, editor. Operation Epsilon: the Farm Hall transcripts. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1993. ISBN 0-520-08499-3. ix + 313 pp. LCCN Q141 .O56 1993b.

Frayn:19xx:PC

[Fraxx] Michael Frayn. Postscript to Copenhagen. Web document.,

19xx. URL http://www-personal.umich.edu/~sanders/214/
other/news/CopenhagenPostscript.html.

Fierz:1960:TPT

[FW60] Markus Fierz and Victor Frederick Weisskopf. Theoretical physics in the Twentieth Century: a memorial volume to Wolfgang Pauli. Interscience Publishers, New York, NY, USA, 1960. x + 328 pp. LCCN QC3 .F52. See [Gou61c] for comments on some inaccuracies in this book, and on the history of the discovery of the spin of the electron and the nucleus.

Gawlik:1991:HIV

[Gaw91] Wojciech Gawlik. Hyperfine interaction versus strong laser fieldoptical Back–Goudsmit (Paschen–Back) effect. American Journal of Physics, 59(8):706–710, August 1991. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Goudsmit:1927:KQN

[GB27a] S. Goudsmit and E. Back. Die Koppelung der Quantenvektoren bei Neon, Argon und einigen Spektren der Kohlenstoffgruppe. (German) [Coupling of quantum vectors in neon, argon and some carbon group spectra]. Zeitschrift für Physik, 40(7):530–538, July 1927. CODEN ZEPYAA. ISSN 0044-3328. URL http://link.

springer.com/article/10.1007/BF01440830.

Goudsmit:1927:FTW

[GB27b] S. Goudsmit and E. Back. Feinstrukturen und Termordnung des Wismutspektrums. (German) [Fine structure and term order of the bismuth spectrum]. Zeitschrift für Physik, 43(5–6): 321–334, May 1927. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF01397446;   http://www.springerlink.
com/content/h5w4160831u31v16/.

Goudsmit:1929:PBE

[GB29a] S. Goudsmit and R. F. Bacher. The Paschen–Back effect of hyperfine structure. Physical Review (2), 34(11):1499–1500, December 1929. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRev.34.1499;   http://prola.aps.org/abstract/
PR/v34/i11/p1499_1.

Goudsmit:1929:SHS

[GB29b] S. Goudsmit and R. F. Bacher. Separations in hyperfine structure. Physical Review (2), 34(12):1501–1506, December 1929. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1930:PBE

[GB30] S. Goudsmit and R. F. Bacher. Der Paschen–Back-Effekt der Hyperfeinstruktur. (German) [The Paschen–Back effect in hyperfine structure]. Zeitschrift für Physik, 66(1–2):13–30, January 1930. CODEN ZEPYAA. ISSN 0044-3328. URL http://link.

springer.com/article/10.1007/BF01397522.

Goudsmit:1933:AHS

[GB33] S. Goudsmit and R. F. Bacher. Anomalies in hyperfine structure. Physical Review (2), 43(11):894–899, June 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1967:ZGT

[GC67] Samuel A. (Samuel Abraham) Goudsmit and Robert Claiborne. Die Zeit. (German) [Time]. Time-Life International, Amsterdam, The Netherlands, 1967. 200 pp.

Goudsmit:1969:LTF

[GC69] Samuel A. (Samuel Abraham) Goudsmit and Robert Claiborne. L’homme et le temps. (French) [Man and time]. Time-Life, New York, NY, USA, 1969. 200 pp. Translation from English to French by Serge Ouvaroff.

Goudsmit:1970:MDT

[GC70a] Samuel A. Goudsmit and Robert Clairborne. La mesure du temps. (French) [The measure of time], volume 16 of Sciences. Laffont, Paris, France, 1970. 190 pp. Translation from English to French by Serge Ouvaroff.

Goudsmit:1970:ZGT

[GC70b] Samuel A. (Samuel Abraham) Goudsmit and Robert Claiborne. Die Zeit. (German) [Time]. Rowohlt, Reinbek bei Hamburg, Germany, 1970. ISBN 3-499-18016-2 (paperback). 200 (est.) pp.

Goudsmit:1966:T

[GCt66] Samuel A. Goudsmit, Robert Claiborne, and the Editors of Life. Time. Time, Inc., New York, NY, USA, 1966. 200 pp.

Goudsmit:1967:TDT

[GCt67] Samuel A. Goudsmit, Robert Claiborne, and the Editors of Life. De tijd. (Dutch) [Time]. Het Parool, Amsterdam, The Netherlands, 1967. 200 pp. Translation to Dutch by Ph. Polderman-van Gogh from the English original.

Goudsmit:1925:IZCb

[GdLK25a] S. Goudsmit and R. de L. Kronig. De intensiteiten der Zeemancomponenten. (Dutch) [The intensities of the Zeeman components]. Koninklijke Akademie van Wetenschappen te Amsterdam,

24(2):278–281, ???? 1925. URL https://repository.aip.org/
islandora/object/nbla%3A227304#page/6/.

Goudsmit:1925:IZG

[GdLK25b] S. Goudsmit and R. de L. Kronig. Die Intensität der Zeemankomponenten. (German) [The intensity of the Zeeman components]. Die Naturwissenschaften, 13(5):90, January 1925. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http:

//link.springer.com/article/10.1007/BF01558737.

Goudsmit:1925:IZCa

[GdLK25c] S. Goudsmit and R. de L. Kronig. The intensities of the Zeeman components. Koninklijke Akademie van Wetenschappen te Amsterdam, 28(4):418–422, ???? 1925. URL https://repository.

aip.org/islandora/object/nbla%3A227304#page/1/.

Goudsmit:1925:U

[GdLK25d] S. Goudsmit and R. de L. Kronig. [unknown]. Verhandelingen — Koninklijke Nederlandse Akademie van Wetenschappen, Afdeling Natuurkunde, Tweede Reeks, 34(??):278–??, ???? 1925. CODEN VNAWAG. ISSN 0373-465X.

Goudsmit:1958:Ec

[GdMMP58] S. A. Goudsmit, E. der Mateosian, M. McKeown, and S. Pasternack. Editorial. Physical Review Letters, 1(6):197–198, September 15, 1958. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1944:SFN

[GF44] S. A. Goudsmit and W. H. Furry. Significant figures of numbers in statistical tables. Nature, 154(3921):800–801, December 23, 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v154/

n3921/pdf/154800a0.pdf. See later work [LG78], and severe criticism [Rai76, Rai85].

Goudsmit:1931:MES

[GG31] S. Goudsmit and L. Gropper. Many-electron selection rules. Physical Review (2), 38(2):225–236, July 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1928:MS

[GH28] S. Goudsmit and C. J. Humphreys. Multiplet separations. Physical Review (2), 31(6):960–966, June 1928. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1931:HSI

[GI31] S. Goudsmit and D. R. Inglis. The hyperfine structure of ionized lithium. Physical Review (2), 37(3):328–329, February 1931. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Giulini:2008:ESC

[Giu08] Domenico Giulini. Electron spin or “classically non-describable two-valuedness”. Studies in History and Philosophy of Modern Physics, 39(3):557–578, September 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219808000269.

Glanz:2000:PFA

[Gla00] James Glanz. Of physics, friendship and the atomic bomb: 60 years later, a mystery abides of physics, friendship and Nazi Germany’s atomic bomb efforts mystery of a 1941 meeting becomes drama on the stage. New York Times, ??(??):F1–F2, March 21, 2000. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/91737705/.

Goudsmit:1933:VLO

[GO33] S. Goudsmit and J. S. Owens. Viscosity of lubricating oils. Submitted to Industrial and Engineering Chemistry, but rejected on 26 October 1933 because of subject overlap with an already-published

paper., 1933. URL https://repository.aip.org/islandora/
object/nbla:229151#page/40/mode/1up.

Goldhaber:1979:OSG

[Gol79] Maurice Goldhaber. Obituary: Samuel A. Goudsmit. Physics Today, 32(4):71–72, April 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.2995511.

Goldberg:1989:BON

[Gol89] Stanley Goldberg. Between old and new: Goudsmit at Brookhaven. In De Maria et al. [DGS89], pages 125–129. ISBN 9971-5-0740-4. LCCN QC9.U5 I57 1988.

Goudsmit:1921:RAD

[Gou21] S. Goudsmit. Relativistische Auffassung des Dubletts. (German) [Relativistic view of the doublet]. Die Naturwissenschaften, 9 (49):995, December 1921. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://link.springer.

com/article/10.1007/BF01489623.

Goudsmit:1922:DSV

[Gou22] S. Goudsmit. Les doublets dans les spectres visibles. Archives Neérlandaises des Sciences Exactes et Naturelles, Série III A, 6

(???):116–126, ???? 1922. URL https://repository.aip.org/
islandora/object/nbla%3A226659#page/15/mode/1up.

Goudsmit:1924:ZIS

[Gou24a] S. Goudsmit. Der Zeemaneffekt im Scandiumspektrum. (German) [The Zeeman effect in the scandium spectrum]. Die Naturwissenschaften, 12(37):743–744, September 1924. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http:

//link.springer.com/article/10.1007/BF01504829.

Goudsmit:1924:SI

[Gou24b] S. Goudsmit. The spectrum of iron. Nature, 113(2843):604, April 26, 1924. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Goudsmit:1924:SM

[Gou24c] S. Goudsmit. The spectrum of manganese. Nature, 113(2833):238, February 16, 1924. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Goudsmit:1924:OLG

[Gou24d] S. Goudsmit. Zur Ordnung des Lanthanspektrums. (German) [On the ordering of the lanthanum spectrum]. Die Naturwissenschaften, 12(41):851–852, October 1924. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://

link.springer.com/article/10.1007/BF01507551.

Goudsmit:1925:BVL

[Gou25a] S. Goudsmit. De bouw van het lanthanspectrum. (Dutch) [The construction of the lanthanum spectrum]. In Handelingen van het XXe Nederlandsch Natuur- en Geneeskungdig Congres, gehouden op 14, 15 en 16 April 1925 te Groningen, pages 1–2. ????, ????,

1925. URL https://repository.aip.org/islandora/object/
nbla%3A227110#page/1.

Goudsmit:1925:DTL

[Gou25b] S. Goudsmit. Determinations of the terms in the lanthanum spectrum. Koninklijke Akademie van Wetenschappen te Amsterdam,

28(1):23–31, ???? 1925. URL https://repository.aip.org/
islandora/object/nbla%3A227114#page/1/.

Goudsmit:1925:GND

[Gou25c] S. Goudsmit. Het grondniveau in het neonspectrum. (Dutch) [The ground state in the neon spectrum]. Physica (The Hague),

5(??):70–71, ???? 1925. URL https://repository.aip.org/
islandora/object/nbla%3A227267#page/1.

Goudsmit:1925:ISA

[Gou25d] S. Goudsmit. Iets over spectra en atoombouw. (Dutch) [Something about spectra and atomic construction]. Physica (The Hague), 5(??):281–292, ???? 1925. CODEN PYSIA7. ISSN 0370-

2707. URL https://repository.aip.org/islandora/object/
nbla%3A227272#page/1/.

Goudsmit:1925:TLS

[Gou25e] S. Goudsmit. Termbepalingen in het lanthaan-spectrum. (Dutch) [Determinations of the terms in the lanthanum spectrum]. Koninklijke Akademie van Wetenschappen te Amsterdam, 33(8):774–782

+ 226–229, ???? 1925. URL https://repository.aip.org/
islandora/object/nbla%3A227114#page/11/.

Goudsmit:1925:GNG

[Gou25f] S. Goudsmit. Über den Grundterm des Neonspektrums. (German) [On the basic term of the neon spectrum]. Zeitschrift für Physik, 32(1):111–112, December 1925. CODEN ZEPYAA. ISSN 0044-

3328. URL http://link.springer.com/article/10.1007/
BF01331651;   https://repository.aip.org/islandora/object/
nbla%3A227365#page/1.

Goudsmit:1925:WTS

[Gou25g] S. Goudsmit. Über die g-Werte der Terme in Spektren höherer Stufe. (German) [On the g values of the terms in higherlevel spectra]. Die Naturwissenschaften, 13(51):1090, December 1925. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic).  URL http://link.springer.com/
article/10.1007/BF01585593;   https://repository.aip.
org/islandora/object/nbla%3A227391#page/1/.

Goudsmit:1925:KSG

[Gou25h] S. Goudsmit. Über die Komplexstruktur der Spektren. (German) [On the complex structure of spectra]. Zeitschrift für Physik, 32 (1):794–798, December 1925. CODEN ZEPYAA. ISSN 0044-

3328. URL http://link.springer.com/article/10.1007/
BF01331715;   https://repository.aip.org/islandora/object/
nbla%3A227609#page/3/.

Goudsmit:1927:ASS

[Gou27] Samuel Abraham Goudsmit. Atoommodel en structuur der spectra. (Dutch) [Atomic Model and Structure of Spectra]. Ph.D. thesis, Universiteit Leiden, Leiden, The Netherlands, 1927. x + 68

pp. URL https://genealogy.math.ndsu.nodak.edu/id.php?
id=75194.

Goudsmit:1928:MSE

[Gou28] S. Goudsmit. Multiplet separations for equivalent electrons and the Roentgen doublet law. Physical Review (2), 31(6):946–959, June 1928. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1929:GII

[Gou29a] S. Goudsmit. Gibt es eine Isotopenverschiebung im Cadmiumspektrum?. (German) [Is there an isotope shift in the cadmium spectrum?]. Die Naturwissenschaften, 17(41):805–806, October 1929.

CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01506659.

Goudsmit:1929:QM

[Gou29b] S. A. Goudsmit. Quantum mechanics. Journal of The Franklin Institute, 207(4):523–524, 1929. CODEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0016003229918354.

Goudsmit:1930:EHS

[Gou30a] S. Goudsmit. An extension of Houston’s and Slater’s multiplet relations. Physical Review (2), 35(11):1325–1336, June 1930. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1930:PMH

[Gou30b] S. Goudsmit. Predictions of manganese hyperfine structure. Physical Review (2), 35(4):440, February 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1930:RHS

[Gou30c] S. Goudsmit. A remark on hyperfine structure. Physical Review (2), 35(4):436–437, February 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1931:THS

[Gou31] S. Goudsmit. Theory of hyperfine structure separations. Physical Review (2), 37(6):663–681, March 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1932:PDT

[Gou32] Samuel Goudsmit. Present difficulties in the theory of hyperfine structure. In Convegno di Fisica Nucleare, Ottobre 1931, pages 33–49. Reale Accademia d’Italia, Rome, Italy, 1932.

Goudsmit:1933:NMM

[Gou33] S. Goudsmit. Nuclear magnetic moments. Physical Review (2), 43(8):636–639, April 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1936:SN

[Gou36] S. Goudsmit. On the slowing down of neutrons. Physical Review (2), 49(5):406, March 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1937:SS

[Gou37] S. Goudsmit. Symmetry of symbols. Nature, 139(3514):417, March 6, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Goudsmit:1939:RBE

[Gou39] S. Goudsmit. Radioactivity in biological experiments. Science, 90(2348):615–616, December 29, 1939. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Goudsmit:1942:RCI

[Gou42] Samuel A. Goudsmit. Reflection coefficients and impedance charts. Report, Radiation Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1942. 21 pp. URL https://lib.

mit.edu/record/cat00916a/mit.000235861.

Goudsmit:1943:CBS

[Gou43] Samuel A. Goudsmit. The comparison between signal and noise. Report, Radiation Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1943. 6 pp.

Goudsmit:1945:RDL

[Gou45] S. Goudsmit. Random distribution of lines in a plane. Reviews of Modern Physics, 17(2–3):321–322, April 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.17.321;
 http://rmp.aps.org/abstract/RMP/v17/i2-3/p321_1.

Goudsmit:1946:NCC

[Gou46a] S. A. Goudsmit. The Nazis came close — but. Chemical & Engineering News, 24(16):2176, August 1946. CODEN CENEAR. ISSN 0009-2347 (print), 1520-605X (electronic).

Goudsmit:1946:HGL

[Gou46b] Samuel A. Goudsmit. How Germany lost the race. Bulletin of the Atomic Scientists, 1(7):4–5, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1946:WPG

[Gou46c] Samuel A. Goudsmit. War physics in Germany. Review of Scientific Instruments, 17(??):49–52, ???? 1946. CODEN RSINAK. ISSN 1089-7623, 0034-6748.

Goudsmit:1946:SS

[Gou46d] Samuel Abraham Goudsmit. Secrecy or science? Science Illustrated, 1(??):6–9, May 1946.

Goudsmit:1947:HAC

[Gou47a] S. Goudsmit. Heat of atomization of carbon. Nature, 159(4048): 742–743, May 31, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Goudsmit:1947:GAR

[Gou47b] Samuel A. Goudsmit. German atom research: Scientist disputes statement on development of bomb: the writer of the following letter, Professor of Physics at Northwestern University, headed the American mission investigating Germany’s progress in atomic research. New York Times, ??(??):E8, November 9, 1947. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/108045246/.

Goudsmit:1947:GSA

[Gou47c] Samuel A. Goudsmit. German scientists in Army employment: I. The case analyzed. Bulletin of the Atomic Scientists, 3(2):64, 67, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1947:HGN

[Gou47d] Samuel A. Goudsmit. Heisenberg on the German Nuclear Power Project. Bulletin of the Atomic Scientists, 3(11):343, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Response to [Hei47a].

Goudsmit:1947:PA

[Gou47e] Samuel A. Goudsmit. [preview of Alsos]. Life Magazine, page ??, October 20, 1947.

Goudsmit:1947:SAB

[Gou47f] Samuel A. Goudsmit. Science and the atom bomb: Secret’s discovery by other nations is held inevitable. New York Times, ??

(??):26, February 11, 1947. URL http://search.proquest.com/
hnpnewyorktimes/docview/107824870/.

Goudsmit:1947:A

[Gou47g] Samuel Abraham Goudsmit. Alsos. H. Schuman, New York, NY, USA, 1947. xiv + 259 pp. LCCN D810.S2 G6 1947.

Goudsmit:1947:AFG

[Gou47h] Samuel Abraham Goudsmit. Alsos: the failure in German science. Sigma Books, London, UK, 1947. xiv + 259 pp. LCCN D810.S2 G6 1947a.

Goudsmit:1947:NAS

[Gou47i] Samuel Abraham Goudsmit. Nazis’ atomic secrets. Life, 23(??): 123, October 1947.

Goudsmit:1948:TFM

[Gou48a] S. A. Goudsmit. A time-of-flight mass spectrometer. Physical Review (2), 74(5):622–623, September 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1948:OTG

[Gou48b] Samuel A. Goudsmit. Our task in Germany. Bulletin of the Atomic Scientists, 4(4):106, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1948:LSA

[Gou48c] Samuel Abraham Goudsmit. L’Allemagne et le secret atomique (la mission Alsos). (French) [Germany and the atomic secret (Alsos mission)]. A. Fayard, Paris, France, 1948. 250 + 1 pp. LCCN D810.S2 G615.

Goudsmit:1949:GWR

[Gou49a] Samuel A. Goudsmit. German war research: Scientist’s claim of progress with atom is disputed. New York Times, ??(??): E8, January 9, 1949. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/105811389/. The writer of the

following letter, co-discovered of the “spin of the Electron”, was detailed to the War Department as chief of the ALSOS mission investigating Germany’s progress in Atomic research. He is now a senior physicist at the Brookhaven National Laboratory.

Goudsmit:1949:SAK

[Gou49b] Samuel A. Goudsmit. Soviet atom knowledge. New York Times, ?? (??):E8, December 18, 1949. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/105885035/.

Goudsmit:1950:BRJ

[Gou50a] S. A. Goudsmit. Book review: Journal of Applied Mathematics and Physics (Zeitschrift für Angewandte Mathematik und Physik), Vol. 1, No. 1, January 15, 1950. Edited by R. Sänger, Verlag Birkhäuser, Basel, Switzerland. Physics Today, 3(4):35–36, April 1, 1950. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL https://physicstoday.scitation.org/doi/
abs/10.1063/1.3066867.

Goudsmit:1950:FS

[Gou50b] S. A. Goudsmit. Freedom of science. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 94(2):111–113, 1950. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.

org/stable/3143209.

Goudsmit:1951:A

[Gou51] Samuel A. Goudsmit. Alsos. Sigma Books, London, UK, 1951. ???? pp.

Goudsmit:1953:BRC

[Gou53a] S. A. Goudsmit. Book review: Crime Investigation, by Paul L. Kirk. 765 pp. Interscience Publishers, Inc. New York, 1953. Physics Today, 6(5):17, May 1, 1953. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.3061228.

Goudsmit:1953:BRM

[Gou53b] S. A. Goudsmit. Book review: Marx and Science, by J. D. Bernal. 48 pp. International Publishers, New York, 1952. Physics Today, 6(6):14–15, June 1, 1953. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://physicstoday.

scitation.org/doi/abs/10.1063/1.3061256.

Goudsmit:1955:TSO

[Gou55] S. A. Goudsmit. The task of the security officer. Bulletin of the Atomic Scientists, 11(4):145–147, April 1955. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL https://

lib.mit.edu/record/rgr/523070848. Special issue on Secrecy,

Security, and Loyalty.

Goudsmit:1956:SI

[Gou56] Samuel A. Goudsmit. Scientific intelligence. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 100(1):100–103, 1956. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://

www.jstor.org/stable/3143810.

Goudsmit:1958:Eb

[Gou58a] S. A. Goudsmit. Editorial. Physical Review Letters, 1(5):165–166, September 1, 1958. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1958:Ed

[Gou58b] S. A. Goudsmit. Editorial. Physical Review Letters, 1(12):437–438, December 15, 1958. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1958:IAa

[Gou58c] S. A. Goudsmit. Important announcement. Physical Review (2), 110(1):292, April 1958. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1958:IAb

[Gou58d] S. A. Goudsmit. Important announcement. Physical Review (2), 110(2):582, April 15, 1958. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1958:IAc

[Gou58e] S. A. Goudsmit. Important announcement. Physical Review (2), 110(3):767, May 1958. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1958:LE

[Gou58f] S. A. Goudsmit. A letter from an Editor. Physics Today, 11(11): 55–??, November 1, 1958. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://physicstoday.

scitation.org/doi/abs/10.1063/1.3062299.

Goudsmit:1958:FKP

[Gou58g] Samuel A. Goudsmit, editor. Frontiers of knowledge: papers read at the inauguration of the Graduate School of Arts and Sciences of Brandeis University. Brandeis University, Waltham, MA, USA, 1958. 48 pp.

Goudsmit:1958:PE

[Gou58h] Samuel A. Goudsmit. Physics and experience. In Frontiers of knowledge: papers read at the inauguration of the Graduate School of Arts and Sciences of Brandeis University [Gou58g], page ??

Goudsmit:1959:Eb

[Gou59a] S. A. Goudsmit. Editorial. Physical Review Letters, 3(11):505, December 1, 1959. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1959:Ed

[Gou59b] S. A. Goudsmit. Editorial. Physical Review Letters, 3(11):505, December 1, 1959. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1959:FB

[Gou59c] S. A. Goudsmit. The fermi and the barn. Physical Review Letters, 3(4):161–162, August 15, 1959. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1959:LEF

[Gou59d] S. A. Goudsmit. Letter to the Editor: Fallout fallacies. Bulletin of the Atomic Scientists, 15(7):321, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1959:M

[Gou59e] S. A. Goudsmit. Masers. Physical Review Letters, 3(3):125, August 1, 1959. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1959:N

[Gou59f] S. A. Goudsmit. Nomenclature. Physical Review Letters, 3(5): 205–206, September 1, 1959. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1959:W

[Gou59g] S. A. Goudsmit. A warning. Physical Review Letters, 2(3):79–80, February 1, 1959. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1960:Ea

[Gou60a] S. A. Goudsmit. Editorial. Physical Review Letters, 4(7):335, April 1, 1960. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1960:Eb

[Gou60b] S. A. Goudsmit. Editorial. Physical Review Letters, 4(8):395– 396, April 15, 1960. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1960:Ec

[Gou60c] S. A. Goudsmit. Editorial. Physical Review Letters, 5(6):233, September 15, 1960. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1960:P

[Gou60d] S. A. Goudsmit. Publicity. Physical Review Letters, 4(1):1–2, January 1, 1960. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1961:E

[Gou61a] S. A. Goudsmit. Editorial. Physical Review Letters, 6(11):587– 588, June 1, 1961. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1961:LEB

[Gou61b] S. A. Goudsmit. Letter to the Editor: Big risks and small. Bulletin of the Atomic Scientists, 17(9):393, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1961:PNS

[Gou61c] S. A. Goudsmit. Pauli and nuclear spin. Physics Today, 14 (6):18–21, June 1, 1961. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v14/i6/p18/s1;   https://physicstoday.
scitation.org/doi/abs/10.1063/1.3057597. See [FW60] for

the book whose inaccuracies are addressed in this article, which also records the author’s involvement in the historical discovery of the spin of the electron and the nucleus.

Goudsmit:1962:MA

[Gou62] Samuel A. Goudsmit. Missiia Alsos. ????, ????, 1962. ??–?? pp.

Goudsmit:1963:LEC

[Gou63] S. A. Goudsmit. Letter to the Editor: “credo of an angry old man”. Bulletin of the Atomic Scientists, 19(5):33–34, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1965:BRP

[Gou65a] S. A. Goudsmit. Book review: Proper Words in Proper Places, by Irving T. Richards and Paul I. Richards. 206 pp. The Christopher Publishing House. Boston, Mass., 1964. Physics Today, 18 (3):70, 72, March 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://physicstoday.

scitation.org/doi/abs/10.1063/1.3047281.

Goudsmit:1965:CSL

[Gou65b] S. A. Goudsmit. Claim-staking letters. Physical Review Letters, 15(12):505–506, September 20, 1965. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1965:CP

[Gou65c] S. A. Goudsmit. Communication problems. Physical Review Letters, 15(13):543–544, September 27, 1965. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1965:F

[Gou65d] S. A. Goudsmit. Festschriften. Physical Review Letters, 15(6): 229–230, August 9, 1965. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1966:SR

[Gou66a] S. A. Goudsmit. Selecting a referee. Physical Review Letters, 17 (10):507–508, September 5, 1966. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1966:VRP

[Gou66b] S. A. Goudsmit. The value of refereeing papers. Physical Review Letters, 17(16):845–846, October 17, 1966. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1966:LWR

[Gou66c] Samuel A. Goudsmit. Is the literature worth retrieving? Physics Today, 19(9):52–55, September 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.3048467.

Goudsmit:1967:C

[Gou67a] S. A. Goudsmit. “Comments”. Physical Review Letters, 18(9):301, February 17, 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1967:DDS

[Gou67b] S. A. Goudsmit. La découverte du spin de l’électron. (French) [The discovery of electron spin]. Journal de physique (Paris), 28 (1):123–128, 1967. CODEN JOPQAG. ISSN 0302-0738.

Goudsmit:1967:LPR

[Gou67c] S. A. Goudsmit. Let’s not pay referees. Physics Today, 20(11): 9–??, November 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://physicstoday.

scitation.org/doi/abs/10.1063/1.3034046.

Goudsmit:1967:S

[Gou67d] S. A. Goudsmit. Superficiality. Physical Review Letters, 18(6): 193–194, February 6, 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1967:UPN

[Gou67e] S. A. Goudsmit. Unusual prime number sequences. Nature, 214 (5093):1164, June 10, 1967. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Goudsmit:1968:PCPa

[Gou68a] S. Goudsmit. Page-charge problem. Physical Review (2), 175(4): 1233–1234, November 1968. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1968:BJ

[Gou68b] S. A. Goudsmit. A balanced journal. Physical Review Letters, 21(21):1469–1470, November 18, 1968. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1968:EWN

[Gou68c] S. A. Goudsmit. Editorial: Why the new format? Physical Review (2), 175(1):1, November 1968. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1968:IAP

[Gou68d] S. A. Goudsmit. Important announcement: Page-charge problem. Physical Review (2), 176(2):425–426, December 1968. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1968:PCPb

[Gou68e] S. A. Goudsmit. Page-charge problem. Physical Review (2), 175 (5):1553–1554, November 1968. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1968:PCPc

[Gou68f] S. A. Goudsmit. Page charge problem. Physical Review (2), 176 (1):1–2, December 1968. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1968:PCPd

[Gou68g] S. A. Goudsmit. Page-charge problem. Physical Review (2), 175 (3):697–698, November 1968. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1969:PRA

[Gou69a] S. A. Goudsmit. Physical Review abstracts. Physical Review Letters, 23(24):1367, December 15, 1969. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1969:WHM

[Gou69b] Samuel A. Goudsmit. What happened to my paper? Physics Today, 22(5):23–25, May 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.3035571.

Goudsmit:1970:B

[Gou70a] S. A. Goudsmit. Bias. Physical Review Letters, 25(7):419–420, August 17, 1970. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1970:PCa

[Gou70b] S. A. Goudsmit. Page charges. Physical Review A (Atomic, Molecular, and Optical Physics), 2(2):271, August 1970. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Goudsmit:1970:PCb

[Gou70c] S. A. Goudsmit. Page charges. Physical Review B: Solid State, 2 (3):567, August 1970. CODEN PLRBAQ. ISSN 0556-2805.

Goudsmit:1970:PCc

[Gou70d] S. A. Goudsmit. Page charges. Physical Review B: Solid State, 2 (4):785, August 1970. CODEN PLRBAQ. ISSN 0556-2805.

Goudsmit:1970:PCd

[Gou70e] S. A. Goudsmit. Page charges. Physical Review C (Nuclear Physics), 2(2):329, August 1970. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

Goudsmit:1970:PCe

[Gou70f] S. A. Goudsmit. Page charges. Physical Review D (Particles and Fields), 2(3):415, August 1970. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

Goudsmit:1970:PCf

[Gou70g] S. A. Goudsmit. Page charges. Physical Review D (Particles and Fields), 2(4):631, August 1970. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

Goudsmit:1971:OVE

[Gou71a] S. A. Goudsmit. De ontdekking van de electronenrotatie. (Dutch) [The discovery of electron spin]. Nederlands Tijdschrift voor Natuurkunde, 37(16):386–389, ???? 1971. CODEN NTINEL. ISSN

0926-4264. URL https://www.ntvn.nl/archief/pdf/340/4/.

See also English translation [Gou72d].

Goudsmit:1971:PP

[Gou71b] S. A. Goudsmit. Physics personified. Nature Physical Science, 229 (1):31, January 1971. CODEN NPSCA6. ISSN 0300-8746 (print), 2058-1106 (electronic).

Goudsmit:1971:RPCa

[Gou71c] S. A. Goudsmit. Reduction in page charges. Physical Review C (Nuclear Physics), 3(4):1439, April 1971. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

Goudsmit:1971:RPCc

[Gou71d] S. A. Goudsmit. Reduction in page charges. Physical Review C (Nuclear Physics), 3(5):1727, May 1971. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

Goudsmit:1971:DES

[Gou71e] Samuel A. Goudsmit. The discovery of the electron spin. Web

document, April 1971. URL http://rowdysites.msudenver.
edu/~dovej/courses/am/supplemental/goudsmit-spin.pdf;
 http://www.lorentz.leidenuniv.nl/history/spin/goudsmit.

html. English translation by J. H. van der Waals of Goudsmit’s lecture in Dutch for the golden jubilee of the Dutch Physical Society in April 1971.

Goudsmit:1972:AHC

[Gou72a] S. A. Goudsmit. An ad hoc committee. Physical Review Letters, 29(3):143, July 17, 1972. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1972:CAC

[Gou72b] S. A. Goudsmit. Criticism, acceptance criteria, and refereeing. Physical Review Letters, 28(6):331–332, February 7, 1972. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1972:AT

[Gou72c] S. A. Goudsmit. Not for the art trade. Expedition, 14(4):13– 16, Summer 1972. ISSN 0014-4738. URL https://www.penn.

museum/sites/expedition/not-for-the-art-trade/.

Goudsmit:1972:U

[Gou72d] S. A. Goudsmit. [unknown]. Delta, ??(??):77–??, Summer 1972. English translation of [?].

Goudsmit:1972:CAI

[Gou72e] Samuel A. Goudsmit. Cumulative author index of The Physical Review (volumes 141–188) and Physical Review Letters (volumes 16–23) 1966–1969. American Institute of Physics, Woodbury, NY, USA, 1972. ???? pp.

Goudsmit:1973:HWG

[Gou73a] S. A. Goudsmit. How wide is gray? Physical Review Letters, 31(11):669–670, September 10, 1973. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1973:PCP

[Gou73b] S. A. Goudsmit. Publication costs of the Physical Reviews. Physical Review Letters, 31(5):261–262, July 30, 1973. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1973:RPAa

[Gou73c] S. A. Goudsmit. Regarding papers about fundamental theories. Physical Review D (Particles and Fields), 8(2):357, July 1973. CO-DEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

Goudsmit:1973:RPAb

[Gou73d] S. A. Goudsmit. Regarding papers about fundamental theories. Physical Review D (Particles and Fields), 8(3):685–686, August 1973. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

Goudsmit:1973:SA

[Gou73e] S. A. Goudsmit. Secrecy again. Physical Review Letters, 30(1): 1–2, January 1, 1973. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1974:CA

[Gou74a] S. A. Goudsmit. Citation analysis. Science, 183(4120):28, January 11, 1974. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1737230.

Goudsmit:1974:CS

[Gou74b] S. A. Goudsmit. Communications: I am scared. Bulletin of the Atomic Scientists, 30(2):2–3, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The author writes about the failure of the German atomic bomb project in 1944, and American military attitudes about use of atomic weapons. See comment [Sau74].

Goudsmit:1974:IS

[Gou74c] S. A. Goudsmit. An illiterate scribe. American Journal of Archaeology, 78(1):78–79, January 1974. CODEN AJARAE. ISSN 0002-9114 (print), 1939-828X (electronic). URL http://www.jstor.

org/stable/503765.

Goudsmit:1974:PYP

[Gou74d] S. A. Goudsmit. Pay your page charges. Physics Today, 27 (8):9–11, August 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://physicstoday.

scitation.org/doi/abs/10.1063/1.3128794.

Goudsmit:1974:SS

[Gou74e] S. A. Goudsmit. Swan song. Physical Review Letters, 33(17): 991–992, October 21, 1974. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1974:V

[Gou74f] S. A. Goudsmit. Voluntary or not? Science, 183(4126):702, February 22, 1974. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1737646.

Goudsmit:1975:CAI

[Gou75] Samuel A. Goudsmit. Cumulative author index of The Physical Review volumes 81–100, 1951–1955. American Institute of Physics, Woodbury, NY, USA, 1975. ???? pp.

Goudsmit:1976:DMD

[Gou76a] Samuel A. Goudsmit. David M. Dennison. Nature, 262(5565): 240, July 1, 1976. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Goudsmit:1976:FYS

[Gou76b] Samuel A. Goudsmit. Fifty years of spin: It might as well be spin. Physics Today, 29(6):40–43, June 1, 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/www.physicstoday.org/resource/1/phtoad/v29/i6/p40_s1;
 https://physicstoday.scitation.org/doi/abs/10.1063/1.

3023518. See correspondence [RG76].

Goudsmit:1976:ODM

[Gou76c] Samuel A. Goudsmit. Obituary: David M. Dennison. Nature, 262 (5565):240, July 15, 1976. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/262240a0.pdf.

Goudsmit:1976:OGH

[Gou76d] Samuel A. Goudsmit. Obituary: Gustav Hertz. Physics Today, 29 (1):83–??, January 1, 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://physicstoday.

scitation.org/doi/abs/10.1063/1.3023276.

Goudsmit:1976:OSP

[Gou76e] Samuel A. Goudsmit. Obituary: Simon Pasternack. Physics Today, 29(4):87–88, April 1, 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.3023452.

Goudsmit:1976:WH

[Gou76f] Samuel A. Goudsmit. Werner Heisenberg (1901–1976). Yearbook of the American Philosophical Society, pages 74–80,

1976. URL http://www.aip.org/history/nbl/collections/
goudsmit/colls/box46/box46f148.html;   https://repository.
aip.org/islandora/object/nbla:243957#page/1/mode/2up.

Goudsmit:1977:PEP

[Gou77a] Samuel Goudsmit. Pitfalls in elementary probability. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 121(2):188–189, April 29, 1977. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

http://www.jstor.org/stable/986527. Based on a lecture pre-

sented November 11, 1976.

Goudsmit:1977:ID

[Gou77b] Samuel A. Goudsmit. An irrelevant debate. Bulletin of the Atomic Scientists, 33(3):67, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1978:LNG

[Gou78a] S. A. Goudsmit. Letter: Numbers game. Bulletin of the Atomic Scientists, 34(1):6–7, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tho77].

Goudsmit:1978:BSU

[Gou78b] Samuel Goudsmit. Books: Scientists under Hitler: Politics and the Physics Community in the Third Reich, by Alan D. Beyerchen. Bulletin of the Atomic Scientists, 34(7):47–49, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1981:BHF

[Gou81] S. A. Goudsmit. The backview of human figures in ancient Egyptian art. Journal of Near Eastern Studies, 40(1):43–46, 1981. URL

http://www.jstor.org/stable/544421.

Goudsmit:1983:A

[Gou83] Samuel Abraham Goudsmit. Alsos. History of modern physics, 1800–1950. Tomash Publishers, Los Angeles, CA, USA, 1983. ISBN 0-938228-09-9. xxx + 259 pp. LCCN D810.S2 G6 1983.

Goudsmit:1996:A

[Gou96] Samuel Abraham Goudsmit. Alsos. History of modern physics and astronomy. American Institute of Physics, Woodbury, NY, USA, 1996. ISBN 1-56396-415-5. xxxviii + 259 pp. LCCN D810.S2 G6 1996.

Goudsmit:1998:DES

[Gou98] S. A. Goudsmit. The discovery of the electron spin. In Foundations of Modern EPR [EES98], chapter A.1, pages 1– 12. ISBN 981-02-3295-0, 981-281-676-3. LCCN QC763 .E28

1998. URL http://www.worldscientific.com/worldscibooks/

10.1142/3624\#t=toc. Translated by J. H. van der Waals.

Goudsmit:1959:Ea

[GP59a] S. A. Goudsmit and S. Pasternack. Editorial. Physical Review Letters, 2(1):1, January 1, 1959. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1959:Ec

[GP59b] S. A. Goudsmit and S. Pasternack. Editorial. Physical Review Letters, 2(1):1, January 1, 1959. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1959:Ex

[GP59c] S. A. Goudsmit and S. Pasternack. Editorial. Physical Review Letters, 3(1):3, July 1, 1959. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goldberg:1992:DFR

[GP92] Stanley Goldberg and Thomas Powers. Declassified files reopen “Nazi bomb” debate. Bulletin of the Atomic Scientists, 48(7):32– 35, 38–40, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [DR92].

Goudsmit:1964:OES

[GR64] S. A. Goudsmit and Paul I. Richards. The order of electron shells in ionized atoms. Proceedings of the National Academy of Sciences of the United States of America, 51(4):664–671, 1964. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL

http://www.jstor.org/stable/71986.

Goudsmit:1939:MSC

[GS39] S. Goudsmit and J. L. Saunderson. The multiple scattering of charged particles. Physical Review (2), 56(1):122–123, July 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1940:MSEb

[GS40a] S. Goudsmit and J. L. Saunderson. Multiple scattering of electrons. Physical Review (2), 57(1):24–29, January 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1940:MSEc

[GS40b] S. Goudsmit and J. L. Saunderson. Multiple scattering of electrons. Physical Review (2), 57(6):552, March 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1940:MSEa

[GS40c] S. Goudsmit and J. L. Saunderson. Multiple scattering of electrons. II. Physical Review (2), 58(1):36–42, July 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1948:PMM

[GS48] S. A. Goudsmit and W. T. Scott. Proposed method for measuring scattering of particle tracks. Physical Review (2), 74(10): 1537–1538, November 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

George:1966:SELa

[GS66a] G. G. George and P. R. C. Stevens. Scattering and energy loss of electrons in silicon near the radiation damage threshold. Nuclear Instruments and Methods, 43(1):278–284, 1966. CO-DEN NUIMAL. ISSN 0029-554x (print), 1878-3759 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0029554X66905477. Proceedings of the Tenth Summer Meeting of Nuclear Physicists.

George:1966:SELb

[GS66b] G. G. George and P. R. C. Stevens. Scattering and energy loss of electrons in silicon near the radiation damage threshold. Nuclear Instruments and Methods, 43(2):278–284, 1966. CO-DEN NUIMAL. ISSN 0029-554x (print), 1878-3759 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0029554X6690471X.

Goudsmit:1958:Ea

[GT58a] S. A. Goudsmit and George L. Trigg. Editorial. Physical Review Letters, 1(1):1–2, July 1, 1958. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1958:Ee

[GT58b] S. A. Goudsmit and George L. Trigg. Editorial. Physical Review Letters, 1(2):49–50, July 15, 1958. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1958:Ef

[GT58c] S. A. Goudsmit and George L. Trigg. Editorial. Physical Review Letters, 1(3):89–90, August 1, 1958. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1966:AL

[GT66a] S. A. Goudsmit and George L. Trigg. Abstracts for letters. Physical Review Letters, 16(20):885–886, May 16, 1966. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1966:GNR

[GT66b] S. A. Goudsmit and George L. Trigg. Growth and new rules. Physical Review Letters, 17(2):67–68, July 11, 1966. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1966:ROA

[GT66c] S. A. Goudsmit and George L. Trigg. Revised ordering — again. Physical Review Letters, 16(1):1–2, January 3, 1966. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1967:GPR

[GT67a] S. A. Goudsmit and George L. Trigg. Growth of Physical Review Letters. Physical Review Letters, 18(1):1–2, January 2, 1967. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1967:RLL

[GT67b] S. A. Goudsmit and George L. Trigg. Restriction on length of letters. Physical Review Letters, 19(8):413, August 21, 1967. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1968:AL

[GT68] S. A. Goudsmit and George L. Trigg. Abstracts for letters. Physical Review Letters, 20(10):481–482, March 4, 1968. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1969:PC

[GT69] S. A. Goudsmit and George L. Trigg. A policy change. Physical Review Letters, 23(12):629, September 22, 1969. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1970:O

[GT70a] S. A. Goudsmit and George L. Trigg. Oversupply. Physical Review Letters, 25(16):1077–1078, October 19, 1970. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1970:RR

[GT70b] S. A. Goudsmit and George L. Trigg. Referee reports. Physical Review Letters, 24(6):251–252, February 9, 1970. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1970:REO

[GT70c] S. A. Goudsmit and George L. Trigg. Relaxation, in experiment only. Physical Review Letters, 24(2):37–38, January 12, 1970. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1971:TL

[GT71] S. A. Goudsmit and George L. Trigg. Troubles with length. Physical Review Letters, 27(1):1–2, July 5, 1971. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1972:CD

[GT72a] S. A. Goudsmit and George L. Trigg. Critical data. Physical Review Letters, 29(15):973–974, October 9, 1972. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1972:TSE

[GT72b] S. A. Goudsmit and George L. Trigg. Two special errata. A new requirement. Physical Review Letters, 29(26):1717, December 25, 1972. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1974:CL

[GT74] S. A. Goudsmit and George L. Trigg. Comments on letters. Physical Review Letters, 33(4):183–184, July 22, 1974. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Goudsmit:1925:OSV

[GU25] S. Goudsmit and G. E. Uhlenbeck. Opmerking over de spectra van waterstof en helium. (Dutch) [Note on the spectra of hydrogen and helium]. Physica (The Hague), 5(??):266–270, ???? 1925. CODEN

PYSIA7. ISSN 0370-2707. URL https://repository.aip.org/
islandora/object/nbla%3A227343#page/1/.

Goudsmit:1926:KQI

[GU26a] S. Goudsmit and G. E. Uhlenbeck. Die Kopplungsmöglichkeiten der Quantenvektoren im Atom. (German) [The coupling possibilities of quantum vectors in the atom]. Zeitschrift für Physik, 35(8–9):618–625, August 1926. CODEN ZEPYAA.

ISSN 0044-3328. URL http://link.springer.com/article/10.
1007/BF01379808.

Goudsmit:1926:RES

[GU26b] S. Goudsmit and G. E. Uhlenbeck. Over het roteerende electron en de structuur der spectra. (Dutch) [On the rotating electron and the structure of the spectra]. Physica (The Hague), 6(??):273–290, ???? 1926. CODEN PYSIA7. ISSN 0370-2707.

Guerlac:1950:BRA

[Gue50] Henry Guerlac. Book review: Alsos by Samuel A. Goudsmit. Isis, 41(2):224–227, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227208.

Goudsmit:1934:NSC

[GW34] S. Goudsmit and Ta-You Wu. Note on the spectrum of the corona. The Astrophysical Journal, 80(??):154–157, September 1934. CO-DEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

URL https://lib.mit.edu/record/hsr/521369085.

Goudsmit:1943:PMR

[GW43a] Samuel A. Goudsmit and P. R. Weiss. Possible measurement of radar echoes by use of model targets. Report 192, Radiation Laboratory, Massachusetts Institute of Technology, Cambridge,

MA, USA, 1943. 12 pp. URL https://lib.mit.edu/record/
cat00916a/mit.000241321.

Goudsmit:1943:SCN

[GW43b] Samuel A. Goudsmit and P. R. Weiss. Statistics of circuit noise. Technical Report 192, Radiation Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, January 29,

1943. 9 pp. URL https://lib.mit.edu/record/cat00916a/
mit.000241325.

Goudsmit:1930:APR

[GY30] S. Goudsmit and L. A. Young. Atomic physics and related subjects: Communications to Nature: The nuclear moment of lithium. Nature, 125(3151):461–462, March 22, 1930. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Goudsmit:1924:MRS

[GZ24] S. Goudsmit and P. Zeeman. The magnetic resolution of the scandium lines. Nature, 114(2864):432–433, September 20, 1924. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Heisenberg:1927:BGB

[Hei27] Werner Heisenberg. Besprechung: S. Goudsmit und E. Back. Die Kopplung der Quantenvektoren bei Neon, Argon und einigen Spektren der Kohlenstoffgruppe. (German) [Book review: S. Goudsmit und E. Back. Coupling of the quantum vectors in neon, argon, and some spectra of the carbon group]. Physikalische Berichte, 8(??): 470, ???? 1927.

Heisenberg:1946:ATA

[Hei46] Werner Heisenberg. Über die Arbeiten zur Technischen Ausnutzung der Atomkernenergie in Deutschland. (German) [On the work toward the technical use of atomic energy in Germany]. Die Naturwissenschaften, 33(11):325–329, ???? 1946. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http:

//www.springerlink.com/content/j0m0u5j553402765. English

translation in [Hei47a].

Heisenberg:1947:RGTa

[Hei47a] Werner Heisenberg. Research in Germany on the technical application of atomic energy. Nature, 160(4059):211–215, August 16, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v160/

n4059/pdf/160211a0.pdf. Abridged translation of German original [Hei46]. See also [Hei53a, Hei53b], and response [Gou47d].

Heisenberg:1947:RGTb

[Hei47b] Werner Heisenberg. Research in Germany on the technical application of atomic energy. Chemical & Engineering News, 1947 (25):2620–2621, September 15, 1947. CODEN CENEAR. ISSN 0009-2347 (print), 1520-605X (electronic). URL http://pubs.

acs.org/doi/pdf/10.1021/cen-v025n037.p2620.

Heisenberg:1949:GAR

[Hei49] Werner Heisenberg. German atom research: Appraisal of progress said to await release of reports. New York Times, ??(??):8E– 8E, January 30, 1949. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/105774675/. Section 4.

Heisenberg:1953:NPa

[Hei53a] Werner Heisenberg. Nuclear Physics. Methuen and Co. Ltd., London, UK, 1953. ???? pp. LCCN ???? English translation of [HW49].

Heisenberg:1953:NPb

[Hei53b] Werner Heisenberg. Nuclear Physics. Philosophical Library, New York, NY, USA, 1953. ISBN 0-8065-3033-2. ix + 225 pp. LCCN

QC173 .H3854. URL http://alsos.wlu.edu/information.

aspx?id=95. English translation of [HW49]. See also [Hei47a].

Hendricks:1948:LVF

[Hen48] B. Clifford Hendricks. Legend vs. fact: review of Alsos, Samuel A. Goudsmit, xiv + 259 pp., Schuman. New York. 1947. The Scientific Monthly, 66(4):346–347, April 1948. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://

www.jstor.org/stable/19082.

Hoeneveld:2018:BRM

[Hoe18] Friso Hoeneveld. Book review: Martijn van Calmthout. Sam Goudsmit: Zijn jacht op de atoombom van Hitler. Isis, 109(1): 207–208, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hollinger:1993:BRH

[Hol93a] David A. Hollinger. Book review: How uncertain was he?: Heisenberg’s War: The Secret History of the German Bomb. By Thomas Powers. Illustrated 608 pp. New York: Alfred A. Knopf. New York Times, ??(??):BR3–BR4, March 14, 1993. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/109209420/.

Hollinger:1993:EHS

[Hol93b] David A. Hollinger. Eavesdropping on Hitler’s scientists: What 10 German scholars, interned by the Allies in 1945, said when they thought no one was listening. New York Times, ??(??):A32, October 31, 1993. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/109047172/. Review of [Fra93a].

Hays:1951:MMM

[HRG51] E. E. Hays, P. I. Richards, and S. A. Goudsmit. Mass measurements with a magnetic time-of-flight mass spectrometer. Physical Review (2), 84(4):824–829, November 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Hays:1952:MMM

[HRG52] E. E. Hays, P. I. Richards, and S. A. Goudsmit. Mass measurements with a magnetic time-of-flight mass spectrometer. Physical Review (2), 85(6):1065, March 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Howard:2009:RSG

[HS09] William E. Howard III and Lewis E. Snyder. Recollections of Samuel Goudsmit’s humor. Physics Today, 62(10):10, October 1, 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL https://physicstoday.scitation.org/doi/
full/10.1063/1.3248447.

Hermann:1994:CTL

[HUHM94] Gerd Hermann, Thomas Udem, Kichinosuke Hirokawa, and Hideyuki Matsuta. Calculated transmission line structures of a Cu coherent forward scattering resonance monochromator (CF-SRM). Spectrochimica Acta Part B: Atomic Spectroscopy, 49 (4):355–365, 1994. CODEN SAASBH. ISSN 0584-8547 (print),

1873-3565 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0584854794800308.

Heisenberg:1949:PAG

[HW49] Werner Heisenberg and Wilhelm Westphal. Die Physik der Atomkerne. (German) [Physics of the atomic nucleus]. Friedrich Vieweg und Sohn, Braunschweig, Germany, third extended edition, 1949. viii + 192 pp. LCCN ???? English translation in [Hei53a, Hei53b].

Ikegami:2017:EAS

[Ike17] Seiji Ikegami. Extension and applications of switching model: Range theory, multiple scattering model of Goudsmit–Saunderson, and lateral spread treatment of Marwick–Sigmund. Radiation Physics and Chemistry, 138:37–48, 2017. CODEN RPCHDM. ISSN 0969-806X (print), 1879-0895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0969806X17302694.

Ivezic:2010:GUG

[Ive10] Tomislav Ivezić. The generalized Uhlenbeck–Goudsmit hypothesis: ‘magnetic’ S^{a} and ‘electric’ Z^{a} spins. Physica Scripta, 81 (2):025001:1–025001:8, January 2010. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic).

Jewett:1977:AWW

[Jew77] Frank Baldwin Jewett. The Academy in World War II. In The National Academy of Sciences: the first hundred years, 1863–1963 [Coc77], pages 382–432. ISBN 0-309-02518-4. LCCN Q11.N2862

C6. URL https://download.nap.edu/cart/download.cgi?
record_id=579&file=382-432.

Jungk:1958:BTT

[Jun58] Robert Jungk. Brighter Than a Thousand Suns: a Personal History of the Atomic Scientists. Harcourt Brace, New York, NY, USA, 1958. 369 pp. LCCN QC773 .J813. Translated by James Cleugh.

Kaempffert:1947:MKR

[Kae47] Waldemar Kaempffert. Mr. Kaempffert replies [to Samuel A. Goudsmit]. New York Times, ??(??):E8, November 9, 1947. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/108047021/. See [Gou47b].

Kalckar:1985:I

[Kal85] Jørgen Kalckar. Introduction. In Jørgen Kalckar, editor, Foundations of Quantum Physics I (1926–1932), volume 6 of Niels Bohr Collected Works, pages 304–330. Elsevier, Amsterdam, The Netherlands, 1985. ISSN 1876-0503.

URL http://www.sciencedirect.com/science/article/pii/
S1876050308703426.

Karlsch:2005:HBG

[Kar05] Rainer Karlsch. Hitlers Bombe: die geheime Geschichte der deutschen Kernwaffenversuche. (German) [Hitler’s bomb: The Secret History of German Nuclear Tests]. Deutsche Verlags-Anstalt, München, Germany, 2005. ISBN 3-421-05809-1. 415 pp. LCCN

QC773.3.G3 K37 2005. URL http://www.loc.gov/catdir/toc/
fy054/2005433177.html.

Kawrakow:1998:REM

[KB98] Iwan Kawrakow and Alex F. Bielajew. On the representation of electron multiple elastic-scattering distributions for Monte Carlo calculations. Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms, 134(3):325–336, 1998. CODEN NIMBEU. ISSN 0168-583X (print), 1872-9584 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0168583X97007234.

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Kadri:2009:IGS

[KIGT09] O. Kadri, V. Ivanchenko, F. Gharbi, and A. Trabelsi. Incorporation of the Goudsmit–Saunderson electron transport theory in the Geant4 Monte Carlo code. Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms, 267(23):3624–3632, 2009. CO-DEN NIMBEU. ISSN 0168-583X (print), 1872-9584 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0168583X09009823.

Kleppner:2009:WSS

[Kle09] Daniel Kleppner. What Sam said. Physics Today, 62(2):8–9, February 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Kronig:1960:TP

[Kro60] Ralph Kronig. The turning point. In Theoretical physics in the Twentieth Century: a memorial volume to Wolfgang Pauli [FW60], page ?? LCCN QC3 .F52. See [Gou61c] for comments on some inaccuracies in this book, and on the history of the discovery of the spin of the electron and the nucleus.

Kuhn:1963:OHIa

[Kuh63a] Thomas S. Kuhn. Oral history interviews: Samuel A. Goudsmit session I. Web site, December 5, 1963. URL https:/

/www.aip.org/history-programs/niels-bohr-library/oral-
histories/4640-1.

Kuhn:1963:OHIb

[Kuh63b] Thomas S. Kuhn. Oral history interviews: Samuel A. Goudsmit session II. Web site, December 7, 1963. URL https:/

/www.aip.org/history-programs/niels-bohr-library/oral-
histories/4640-2.

Kuhn:1963:OHIc

[Kuh63c] Thomas S. Kuhn. Oral history interviews: Samuel A. Goudsmit session III. Web site, December 7, 1963. URL https:/

/www.aip.org/history-programs/niels-bohr-library/oral-
histories/4640-3.

LaRocque:1978:LSP

[La 78] Gene R. La Rocque. Letter: Strategic parity. Bulletin of the Atomic Scientists, 34(2):3, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tho77].

Lakhtakia:1996:MMH

[Lak96] A. (Akhlesh) Lakhtakia, editor. Models and Modelers of Hydrogen: Thales, Thomson, Rutherford, Bohr, Sommerfeld, Goudsmit, Heisenberg, Schrödinger, Dirac, Sallhofer. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2302-1. xv + 424 pp. LCCN QC171.2 .M63

1996. URL http://www.worldscientific.com/worldscibooks/
10.1142/2788.

Lang:1948:ETA

[Lan48] Daniel Lang. Early Tales of the Atomic Age. Doubleday, Garden City, NY, USA, 1948. 223 pp. LCCN QC173.

Lang:1953:PFSa

[Lan53a] Daniel Lang. Profiles: A farewell to string and sealing wax. I. The New Yorker, ??(??):47–50, 52, 54, 56, 58, 60, 62, 64, 66–68, 70, 72, October 30, 1953. ISSN 0028-

792X. URL https://www.newyorker.com/magazine/1953/10/
30/a-farewell-to-string-and-sealing-wax-1.

Lang:1953:PFSb

[Lan53b] Daniel Lang. Profiles: A farewell to string and sealing wax. II. The New Yorker, ??(??):46–48, 50, 52, 54, 56, 58, 60–63, 66–67, November 14, 1953. ISSN 0028-792X. URL https:

//www.newyorker.com/magazine/1953/11/14/a-farewell-to-
string-and-sealing-wax-2.

Lang:1954:MTL

[Lan54] Daniel Lang. The Man in the Thick Lead Suit: An Account of the Views and Reactions of Various Americans to Current Scientific Problems. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1954. xi + 207 pp. LCCN QC778 .L3166 1954. Collection of the author’s articles on atomic energy in The New Yorker.

Lang:1955:MTL

[Lan55] Daniel Lang. The Man in the Thick Lead Suit: Conversations, Etc., with Atomic Scientists. Gollancz, London, UK, 1955. xi + 207 pp. LCCN QC778 .L3166 1955. Collection of the author’s articles on atomic energy in The New Yorker.

Lang:1959:HMCa

[Lan59a] Daniel Lang. From Hiroshima to the Moon: Chronicles of Life in the Atomic Age. Simon and Schuster, New York, NY, USA, 1959. 496 pp. LCCN QC782 .L3. URL http://books.google.

com/books?id=-jaBAAAAIAAJ.

Lang:1959:HMCb

[Lan59b] Daniel Lang. From Hiroshima to the Moon: Chronicles of Life in the Atomic Age. Dell Publishing, New York, NY, USA, 1959. 543 pp. LCCN QC782 .L3 1959.

Lapp:1959:BRC

[Lap59] Ralph E. Lapp. Book review: Close-ups of scientists at work: From Hiroshima to the Moon: Chronicles of Life in the Atomic Age. By Daniel Lang. 496 pp. New York: Simon & Schuster. New York Times, ??(??):BR3–BR4, November 8, 1959. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/114631640/.

Lasareff:1935:EEA

[Las35] Wladimir Lasareff. États électroniques de l’atome de carbone et calcul des énergies de liaison. (French) [Electronic states of the carbon atom and calculation of the link energies]. Physica, 2 (1):737–742, 1935. CODEN PHYSAG. ISSN 0031-8914 (print),

1873-1767 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0031891435901513.

Laurence:1958:IMW

[Lau58a] William L. Laurence. An indictment of the men who made the bomb: Brighter Than a Thousand Suns. New York Times, ?? (??):BR3, October 12, 1958. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/114521964/. Review of [Jun58].

Laurence:1958:R

[Lau58b] William L. Laurence. A reply. New York Times, ??(??): BR52, November 9, 1958. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/114361108/.

Laurence:1958:SRD

[Lau58c] William L. Laurence. Science in review: Discoveries on the frontiers of physics are explained at a Harvard conference. New York Times, ??(??):E9, March 30, 1958. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/114447661/.

Layzer:1974:RPP

[LG74] Arthur Layzer and S. A. Goudsmit. Referee policies for Physical Review Letters. Physics Today, 27(9):9–??, September 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL https://physicstoday.scitation.org/doi/abs/
10.1063/1.3128850.

Logan:1978:FDP

[LG78] Jonothan L. Logan and Samuel A. Goudsmit. The first digit phenomenon. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 122(4):193–197, August 18, 1978. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://links.jstor.org/sici?
sici=0003-049X(19780818)122:4%3C193:TFDP%3E2.0.CO%3B2-
C;   http://www.jstor.org/stable/986530. This paper con-

tains derivations of both Stigler’s Law and Benford’s Law, and receives strong criticism in [Rai85]. This paper contains an important historical note that is recorded in entry [Ben38].

Lehmann-Haupt:1993:BTD

[LH93] Christopher Lehmann-Haupt. Books of the times: Did a German scientist prevent catastrophe in World War II?: Heisenberg’s War: The Secret History of the German Bomb By Thomas Powers Illustrated. 608 pages. Alfred A. Knopf. New York Times, ?? (??):C13, March 8, 1993. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/109213855/.

Layzer:1964:SOD

[LHLT64] David Layzer, Zdeněk Horák, Margaret N. Lewis, and Daniel P. Thompson. Second-order Z-dependent theory of many-electron atoms. Annals of Physics, 29(1):101–124, 1964. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0003491664901927.

Lang:1954:HCR

[LN54] Daniel Lang and Araújo Netto. O homem com roupa de chumba: (problemas da atualidade). (Portuguese) [The Man in the Thick Lead Suit: current problems]. Livraria Classica Brasileira, Rio de Janeiro, Brazil, 1954. 194 pp. LCCN QC778 .L3166 1954.

Lorentz:1936:U

[Lor36] H. A. Lorentz. [unknown]. In H. A. Lorentz: Collected Works, volume 7, pages 179–?? Nyhoff, The Hague, The Netherlands, 1936.

Lang:1954:MIB

[LR54] Daniel Lang and Hanns C. Retzer. Die Männer im Bleianzug; [The Man in the Thick Lead Suit]. Obelisk Verlag, Vienna, Austria, 1954. 287 pp. LCCN ????

Logan:1991:LEH

[LRD^{+}91] Jonothan L. Logan, Helmut Rechenberg, Max Dresden, A. Van Der Ziel, and Mark Walker. Letters to the Editor: Heisenberg, Goudsmit and the German “A-bomb”. Physics Today, 44(5):13, 15, 90–92, 94–96, May 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/www.physicstoday.org/resource/1/phtoad/v44/i5/p13_s1;
 https://physicstoday.scitation.org/doi/abs/10.1063/1.

2810103. Letters in response to [Wal90], including Walker’s rebuttal.

Levinger:1969:SOE

[LSH^{+}69] J. S. Levinger, Joseph Schaefer, Paul D. Hambourger, S. A. Goudsmit, and George L. Trigg. The second Ovshinsky effect. Physics Today, 22(7):11–??, July 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.3035666.

Luginbill:1969:BRT

[Lug69] George Luginbill. Book review: Time, by Samuel A. Goudsmit, Robert Clairborne, and the Editors of Life, 200 pp., Time, New York, 1987. The American Biology Teacher, 31(4):268, 1969. ISSN 0002-7685 (print), 0002-8444 (electronic), 1938-4211, 1938-422X.

URL http://www.jstor.org/stable/4442548.

MacPherson:1985:STF

[Mac85] Malcolm MacPherson. A small town with a footnote to history: Hidden below, the Nazi A-bomb lab Haigerloch. New York Times, ??(??):10, 47, November 3, 1985. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/111210585.

Morrison:2007:SAW

[Mor07] Margaret Morrison. Spin: All is not what it seems. Studies in History and Philosophy of Modern Physics, 38(3):529–557, September 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000931.

Moyer:1984:BRA

[Moy84] Albert E. Moyer. Book review: Alsos, Samuel A. Goudsmit. The History of Modern Physics, 1800–1950, volume 1. 259 pp. Los Angeles: Tomash, 1983. American Scientist, 72(5):527–528, 1984. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (elec-

tronic). URL http://www.jstor.org/stable/27852953.

Meshkov:1956:ABG

[MU56] S. Meshkov and C. W. Ufford. Application of the Bacher and Goudsmit method to nuclear spectra. Physical Review (2), 101 (2):734–739, January 1956. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Merritt:1933:OPE

[MUG^{+}33] E. Merritt, G. E. Uhlenbeck, S. Goudsmit, G. H. Dieke, and H. W. Stunkard. Obituary: Paul Ehrenfest. Science, 78(2026):377–378, October 27, 1933. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Mosteller:1967:DSR

[MYZ67] Frederick Mosteller, Cleo Youtz, and Douglas Zahn. The distribution of sums of rounded percentages. Demography, 4(2):850– 858, June 1967. CODEN ???? ISSN 0070-3370 (print), 1533-

7790 (electronic). URL http://muse.jhu.edu/journals/dem/
;   http://www.biomedsearch.com/nih/Distribution-Sums-
Rounded-Percentages/21318695.html;   http://www.jstor.

org/stable/2060324. See [DF79] for further work.

Neufeld:1970:AR

[NG70] Jacob Neufeld and S. A. Goudsmit. To amend refereeing. Physics Today, 23(4):9–??, April 1, 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.3022101.

Negreanu:2005:CMS

[NLCS05] C. Negreanu, X. Llovet, R. Chawla, and F. Salvat. Calculation of multiple-scattering angular distributions of electrons and positrons. Radiation Physics and Chemistry, 74(5):264–281, 2005. CODEN RPCHDM. ISSN 0969-806X (print), 1879-0895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0969806X05002239.

Olsen:1963:TGS

[Ols63] Arthur J. Olsen. Trackdown of the German scientist: Nazism and defeat scattered the leaders of a once-great scientific establishment. Herewith a review of where they went and what some of them did when they got there. New York Times, ??(??): 214, September 22, 1963. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/116463805/.

Pais:2000:GSP

[Pai00] Abraham Pais. The Genius of Science: a Portrait Gallery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850614-7 (hardcover). 356 pp. LCCN Q141 .P29 2000.

URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/toc/fy02/99046603.html.

Pash:1969:AM

[Pas69] Colonel Boris T. Pash. The Alsos Mission. Award House, New York, NY, USA, 1969. 256 pp. LCCN D810.S2 P28.

Pash:1980:AM

[Pas80] Colonel Boris T. Pash. The Alsos Mission. Grosset & Dunlap Company, New York, NY, USA, 1980. 256 pp.

Pauli:1924:FTD

[Pau24] W. Pauli. Zur Frage der theoretischen Deutung der Satelliten einiger Spektrallinien und ihrer Beeinflussung durch magnetische Felder. (German) [On the question of the theoretical interpretation of some satellite spectral lines and their influence by magnetic fields]. Die Naturwissenschaften, 12(37):741–743, September 1924. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic).  URL http://link.springer.com/
article/10.1007/BF01504828;   http://www.springerlink.
com/content/q414121432406233/fulltext.pdf. See [GB27b,

BG28] for the first, and independent, interpretation of experimental data for the hyperfine structure of bismuth as being due to a nuclear magnetic moment (spin). See [Gou61c] for corrections to the confusing and often misattributed history of the discovery of [electron and nuclear] spin.

Phelps:1955:TPS

[PBG55] John B. Phelps, Ralph S. Brown, Jr., and S. A. Goudsmit. Toward a positive security program. Bulletin of the Atomic Scientists, 11(4):165–169, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Pfeiffer:1954:BRA

[Pfe54] John Pfeiffer. Book review: Arms and men: The Man in the Thick Lead Suit. By Daniel Lang. 207 pp. New York: Oxford University Press. New York Times, ??(??):BR7, September 12,

1954. URL http://search.proquest.com/hnpnewyorktimes/

docview/113094778/. Collection of the author’s articles on atomic energy in The New Yorker.

Pauling:1930:SLS

[PG30] Linus Pauling and Samuel A. (Samuel Abraham) Goudsmit. The Structure of Line Spectra. International series in physics. McGraw-Hill, New York, NY, USA, 1930. x + 263 pp. LCCN QC451 .P3.

Proca:1939:MPE

[PG39] A. Proca and S. Goudsmit. Sur la masse des particules élémentaires. (French) [On the mass of elementary particles]. Journal de Physique et le Radium, 10(5):209–214, 1939. CODEN JPRAAJ. ISSN 0368-3842.

Pasternack:1961:CAI

[PG61] S. Pasternack and Samuel A. Goudsmit. Cumulative author index of The Physical Review (volumes 101–120) and Physical Review Letters (volumes 1–5) 1951–1955. American Institute of Physics, Woodbury, NY, USA, 1961. ???? pp.

Putte:1974:CGS

[Put74] Dave Walter Vande Putte. Computation of the Goudsmit– Saunderson distribution in electron transport. Nuclear Instruments and Methods, 115(1):117–120, 1974. CODEN NUIMAL. ISSN 0029-554x (print), 1878-3759 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0029554X74904352.

Raimi:1976:FDP

[Rai76] Ralph A. Raimi. The first digit problem. American Mathematical Monthly, 83(7):521–538, August/September 1976. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2319349.

Raimi:1985:FDP

[Rai85] Ralph A. Raimi. The first digit phenomenon again. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 129(2):211–219, June 1985. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL http://www.jstor.org/stable/986989. This paper con-

tains strong criticism of a derivation of Benford’s Law [LG78].

Ruark:1925:CFS

[RC25] Arthur Edward Ruark and Roy L. Chenault. CII. fine structures of spectrum lines. Philosophical Magazine, 50(299):937–956, ???? 1925. CODEN PHMAA4. ISSN 0031-8086.

Rechenberg:1991:HGG

[Rec91] H. Rechenberg. Heisenberg, Goudsmit and the German A-Bomb. Physics Today, 44(5):91–92, May 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v44/i5/p13_s1.

Richtmyer:1976:PP

[RG76] R. D. Richtmyer and Samuel A. Goudsmit. Prejudice in physics? Physics Today, 29(12):42, December 1, 1976. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL https://physicstoday.scitation.org/doi/abs/

10.1063/1.3024639. See [Gou76b, page 42].

Root:1982:SVP

[RG82] Margaret Cool Root and Samuel A. Goudsmit. A scientist views the past: the Samuel A. Goudsmit Collection of Egyptian antiquities; January 30–May 9, 1982, Kelsey Museum of Archaeology. University of Michigan, Ann Arbor, MI, USA, 1982. ii + 44 pp.

Richards:1952:MLB

[RHG52] P. I. Richards, E. E. Hays, and S. A. Goudsmit. Masses of lead and bismuth. Physical Review (2), 85(4):630–631, February 1952. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Ridenour:1947:BRB

[Rid47] Louis N. Ridenour. Book review: Behind the front with a scientific intelligence mission: ALSOS. By Samuel A. Goudsmit. Illustrated. xiv + 459 pp. New York: Henry Schuman. $3.50. New York Times, ??(??):BR8–BR9, November 16, 1947. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/108045114/.

Sauberman:1974:LHS

[Sau74] N. H. Sauberman. Letters: He’s scared, too. Bulletin of the Atomic Scientists, 30(6):6–7, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gou74b].

Segre:1985:HPR

[Seg85] Emilio Segrè. Historical perspective: Refugee scientists and nuclear energy. Annals of the New York Academy of Sciences, 452(1): xv–xix, 1985. CODEN ANYAA9. ISBN 0-89766-298-9, 0-89766-299-7 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). Sixth International Conference on Collective Phenomena: reports from the Moscow Refusnik Seminar / edited by Inga Fischer-Hjalmars and Joel L. Lebowitz. Contributions from the Moscow Refusnik Seminar and from two International Conferences on Collective Phenomena, one held in Stockholm, Sweden, 1–2 December 1983, and the other in Tel Aviv, Israel, 31 May–1 June 1984.

Spitzer:1949:PM

[SHW^{+}49] Lyman Spitzer, Jr., William A. Higinbotham, Arthur S. Wightman, Donald R. Hamilton, David Bohm, Roy Britten, Robert R.

Bush, Elmer G. Butler, Albert Einstein, Luther P. Eisenhart, Samuel A. Goudsmit, M. Stanley Livingston, Stuart Mudd, David Pines, Oswald Veblen, and Irving Wolff. To prevent misunderstanding. Physics Today, 2(8):3, August 1, 1949. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL https://physicstoday.scitation.org/doi/abs/
10.1063/1.3066618.

Sullivan:1978:SGC

[Sul78] Walter Sullivan. Samuel Goudsmit, codiscoverer of electron’s spin, is dead at 76: Many European ties. Discovery at age 11. New York Times, ??(??):B6, December 6, 1978. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/123572175/.

Teng:1986:MCE

[TAL86] S. P. Teng, D. W. Anderson, and D. G. Lindstrom. Monte Carlo electron-transport calculations for clinical beams using energy grouping. International Journal of Radiation Applications and Instrumentation. Part A. Applied Radiation and Isotopes, 37(12):1189–1194, 1986. ISSN 0883-2889 (print), 1878-1128

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0883288986900043.

Temkin:2009:RSG

[Tem09] Aaron Temkin. Recollections of Samuel Goudsmit’s humor. Physics Today, 62(10):10, October 1, 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/full/10.1063/1.4796978.

Thomas:1926:MSE

[Tho26] L. H. (Llewellyn Hilleth) Thomas. The motion of the spinning electron. Nature, 117(2945):514, April 10, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Thompson:1977:LLC

[Tho77] Thurston A. Thompson. Letter: Let’s catch up. Bulletin of the Atomic Scientists, 33(8):2, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Gou78a, La 78, Fel77].

Thomas:1997:FCW

[Tho97] Robert McG. Thomas, Jr. Frederic A. C. Wardenburg 3d, 92, war hero. New York Times, ??(??):44, August 17, 1997. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/109801785/.

Treue:1983:BRP

[Tre83] Wilhelm Treue. Book review: Pride, haughtiness and sobriety among physicists: Alsos, by Samuel A. Goudsmit, with a new introduction by R. V. Jones (Los Angeles/San Francisco: Tomash, 1983), 289 pp. Minerva [Washington, DC], 21(4): 456–459, Winter 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096528.

Torrey:1948:CR

[TWG48] Henry Cutler Torrey, Charles Austin Whitmer, and Samuel Goudsmit, editors. Crystal Rectifiers, volume 15 of Radiation Laboratory Series. McGraw-Hill, New York, NY, USA, 1948. xiii + 443 pp.

Uhlenbeck:1925:EHU

[UG25] George E. Uhlenbeck and Samuel Goudsmit. Ersetzung der Hypothese vom unmechanischen Zwang durch eine Forderung bezüglich des inneren Verhaltens jedes einzelnen Elektrons. (German) [Replacement of the hypothesis of unmechanical force with a requirement for the inner behavior of each electron]. Die Naturwissenschaften, 13(47):953–954, November 1925. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01558878;   http:
//www.springerlink.com/content/u20q85248gp8166q/. See

[Lor36] for a computation in 1927 of the radius of a rotating charged particle with the spin of the electron.

Uhlenbeck:1926:SES

[UG26] G. E. Uhlenbeck and S. Goudsmit. Spinning electrons and the structure of spectra. Nature, 117(2938):264–265, February 20, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). The paper is followed by a half-column supportive letter from Niels Bohr. See [Tho26] for a relativistic correction to the Zeeman splitting of spectral lines.

Uhlenbeck:1929:PBM

[UG29] G. E. Uhlenbeck and S. Goudsmit. A problem in Brownian motion. Physical Review (2), 34(1):145–151, July 1929. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Uhlenbeck:1935:SED

[UG35] G. E. Uhlenbeck and S. Goudsmit. Statistical energy distribution for a small number of particles. In ????, editor, Zeeman Verhandelingen, pages 201–211. Martinus Nijhoff Publishers, Dordrecht, The Netherlands, 1935.

Uhlenbeck:1967:SES

[UG67] G. E. Uhlenbeck and S. Goudsmit. Spinning electrons and the structure of spectra. In W. R. (William Russell) Hindmarsh, editor, Atomic Spectra, The Commonwealth and International Library: Selected Readings in Physics, pages 253–258. Pergamon Press, Oxford, UK and New York, NY, USA, 1967. ISBN 0-08-012103-9, 1-4831-5580-3 (e-book). LCCN QC454 .H626

1967. URL http://www.sciencedirect.com/science/article/

pii/B9780080121031500220. Published in Nature 117, 264 (1926).

Uhlenbeck:1984:SES

[UGB84] G. E. Uhlenbeck, S. Goudsmit, and Niels Bohr. Spinning electrons and the structure of spectra. In Klaus Stolzenburg, editor, The Emergence of Quantum Mechanics (Mainly 1924–1926), volume 5 of Niels Bohr Collected Works, chapter VI, pages 287– 289. Elsevier, Amsterdam, The Netherlands, 1984. ISSN 1876-

0503. URL http://www.sciencedirect.com/science/article/
pii/S1876050308705358.

vanCalmthout:2016:SGJ

[vC16] Martijn van Calmthout. Sam Goudsmit: de jacht op de atoombom van Hitler. Meulenhoff, Amsterdam, The Netherlands, 2016. ISBN 90-290-8958-X (paperback), 94-023-0744-3 (e-book). 284 + 16 pp. LCCN DS135.N6 G656 2016.

vanCalmthout:2018:SGH

[vCH18] Martijn van Calmthout and Michiel Horn, editors. Sam Goudsmit and the hunt for Hitler’s atom bomb. Prometheus Books, Amherst, NY, USA, 2018. ISBN 1-63388-450-3 (hardcover), 1-63388-451-1 (e-book). 243 + 14 pp. LCCN DS135.N6 G65613 2018.

Violino:1972:PCL

[Vio72] Paolo Violino. A program for computing level crossings and the Back–Goudsmit effect. Computer Physics Communications, 4(1):128–137, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900410.

Violino:1984:PCL

[Vio84] Paolo Violino. A program for computing level crossings and the Back–Goudsmit effect. Computer Physics Communications, 35(1–3):C–149–??, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824029.

Walker:1990:HGG

[Wal90] Mark Walker. Heisenberg, Goudsmit and the German Atomic Bomb. Physics Today, 43(1):52–60, January/February 1, 1990. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://alsos.wlu.edu/information.aspx?id=
1572;   http://www.physicstoday.org/resource/1/phtoad/
v43/i1/p52_s1;   https://physicstoday.scitation.org/doi/

abs/10.1063/1.881237. See responses and rebuttal [LRD^{+}91].

Walker:2003:OHV

[Wal03] Mark Walker. Otto Hahn: Verantwortung und Verdrängung. (German) [Otto Hahn: responsibility and repression]. Forschungsprogramm “Geschichte der Kaiser-Wilhelm-Gesellschaft im Nationalsozialismus” (Research Program “History of the Kaiser Wilhelm Society in the National Socialist Era”), Berlin, Germany,

2003. 64 pp. LCCN ???? URL www.mpiwg-berlin.mpg.de/KWG/
Ergebnisse/Ergebnisse10.pdf.

Walker:2006:OHR

[Wal06] Mark Walker. Otto Hahn: Responsibility and repression. Physics in Perspective (PIP), 8(2):116–163, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-006-0277-3.

Werner:2010:ETS

[Wer10] Wolfgang S. M. Werner. Electron transport for spectrum analysis and experiment design. Journal of Electron Spectroscopy and Related Phenomena, 178–179:154–177, 2010. CO-DEN JESRAW. ISSN 0368-2048 (print), 1873-2526 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0368204809002217. Trends in X-ray Photoelectron Spectroscopy of solids (theory, techniques and applications).

Wu:1933:LSH

[WG33] Ta-You Wu and S. Goudsmit. Low states of the heaviest elements. Physical Review (2), 43(6):496, March 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Ward:1967:RAA

[WG67] W. Dixon Ward and S. A. Goudsmit. Reviewer and author anonymity. Physics Today, 20(1):12–??, January 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL https://physicstoday.scitation.org/doi/abs/
10.1063/1.3034118.

Winterberg:1996:MHN

[WHF^{+}96] Friedwardt Winterberg, Günter Herrmann, Igor Fodor, Lincoln Wolfenstein, and Mark E. Singer. More on how Nazi Germany failed to develop the atomic bomb. Physics Today, 49(1):11–83, January 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

White:1948:BRA

[Whi48] Stephen White. Book review: Alsos, by Samuel A. Goudsmit, 259 pp. Henry Schuman, Inc., New York, 1947. Physics Today, 1(3):9, July 1, 1948. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL https://physicstoday.scitation.org/
doi/abs/10.1063/1.3066097.

Wilschut:2017:BRS

[Wil17] H. W. Wilschut. Book review: Sam Goudsmit, by M. van Calmthout. Nederlands Tijdschrift voor Natuurkunde, 83 (3):78–79, March 2017. CODEN NTINEL. ISSN 0926-

4264. URL http://hdl.handle.net/11370/6fb8c57e-1d7d-
4216-9342-e3249be76269;   https://www.ntvn.nl/archief/
pdf/1296/14/;   https://www.rug.nl/research/portal/publications/
sam-goudsmit-by-m-van-calmthout(6fb8c57e-1d7d-4216-9342-
e3249be76269).html.

Winterbon:1987:FAM

[Win87] K. B. Winterbon. Finite-angle multiple scattering. Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms, 21(1):1–7, 1987. CO-DEN NIMBEU. ISSN 0168-583X (print), 1872-9584 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0168583X87901315.

Winterbon:1989:FAM

[Win89] K. B. Winterbon. Finite-angle multiple scattering revisited. Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms, 43(2):146–148, 1989. CODEN NIMBEU. ISSN 0168-583X (print), 1872-9584

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0168583X8990027X.

Zeeman:1930:HWG

[ZBG30] P. Zeeman, E. Back, and S. Goudsmit. Zur Hyperfeinstruktur des Wismuts. (German) [On the hyperfine structure of bismuth]. Zeitschrift für Physik, 66(1–2):1–12, January 1930. CODEN

ZEPYAA. ISSN 0044-3328. URL http://link.springer.com/
article/10.1007/BF01397521;   http://www.springerlink.
com/content/n27vx7850217172r/.

Zahn:1936:MDA

[ZHG36] C. T. Zahn, E. L. Harrington, and S. Goudsmit. A method for the determination of absorption cross sections for thermal neutrons. Physical Review (2), 50(6):570, September 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).